### v1

# A parallel framework for graphical optimal transport

Abstract
We study multi-marginal optimal transport (MOT) problems where the underlying cost has a graphical structure. These graphical multi-marginal optimal transport problems have found applications in several domains including traffic flow control and regression problems in the Wasserstein space. MOT problem can be approached through two aspects: a single big MOT problem, or coupled minor OT problems. In this paper, we focus on the latter approach and demonstrate it has efficiency gain from the parallelization. For tree-structured MOT problems, we introduce a novel parallelizable algorithm that significantly reduces computational complexity. Additionally, we adapt this algorithm for general graphs, employing the modified junction trees to enable parallel updates. Our contributions, validated through numerical experiments, offer new avenues for MOT applications and establish benchmarks in computational efficiency.
11footnotetext: This work was supported by the Swedish Research Council (VR) under grant 2014-5870, SJTU-KTH cooperation grant, the Knut and Alice Wallenberg foundation under grant KAW 2021.0274, and the NSF under grant 2206576.
1 Introduction
Optimal transport (OT) theory offers a fundamental mathematical framework for the comparison of probability distributions (Peyré et al. 2017). Central to OT is the quest for the most cost-efficient strategy to morph one distribution into another. With significant strides in the development of computationally efficient methods (Cuturi 2013, Zanetti and Gondzio 2023), OT has burgeoned into a plethora of applications. It has made substantial impacts in biology (Bunne et al. 2022, Demetci et al. 2020), generative modeling (Fan et al. 2021, 2022b, Korotin et al. 2022), image processing (Peyré et al. 2017, Fan et al. 2021, Zhou and Parno 2024), domain adaptation (Seguy et al. 2017, Perrot et al. 2016), Bayesian statistics (Fan et al. 2020, Srivastava et al. 2015), and control theory (Chen et al. 2016, 2021a, Singh et al. 2022, Ringh et al. 2021a), showcasing the versatility and the expansive scope of OT as a tool for analytical and computational advancement. Multi-marginal optimal transport (MOT) (Ba and Quellmalz 2022, Beier et al. 2023, Trillos et al. 2023, Yang and Zha 2023), a generalization of OT, involves optimizing over transports among more than two distributions, thereby extending the applicability of OT to a broader range of problems (Pass 2012, 2015, Nenna 2016). Its applications are vast and varied, encompassing areas such as signal processing (Elvander et al. 2020), estimation and control (Ringh et al. 2023, 2021a), and machine learning (Haasler et al. 2021c, Noble et al. 2023).
The MOT problem frequently embodies a graphical structure where each node denotes a distribution, and the edges represent the transport plans. For example, the well-known Wasserstein barycenter problem finds a generalized average of a set of distributions (Agueh and Carlier 2011). This is done by finding a distribution that minimizes the sum of optimal transport distances to the given distributions, which can be represented by a star-shaped graph (Fan et al. 2022a). Originally, the Wasserstein barycenter problem was formulated as a sum of coupled bi-marginal OT problems (Carlier and Ekeland 2010, Rabin et al. 2012). However, it can equivalently be posed as a MOT problem, where the graphical structure is instead invoked through the underlying cost. This equivalence between the MOT formulation and the coupled OT formulation holds for any structured problems, where the underlying graph is a tree, i.e., an acyclic graph (Haasler et al. 2021b). Although the coupled OT methodology often presents a more intuitive formulation, its computational development has not been extensively pursued, and existing numerical solvers for general graphical structures rely predominantly on the MOT formulation (Elvander et al. 2020, Haasler et al. 2021b, c, Fan et al. 2022a). To close this gap in the literature, we develop numerical methods for graph-structured OT problems that rely on the coupled OT formulation.
More precisely, following the powerful standard approach of adding an entropic term to the OT objective (Cuturi 2013), we first develop a novel scheme for a sum of OT problems that are coupled according to a tree-structure. We also prove complexity results for solving the problem to a desired accuracy. It turns out that this algorithm can be easily parallelized, thus significantly improving the runtime of previous algorithms that are based on the equivalent MOT problem. We also extend our algorithm to the setting of general graph-structures that may contain cycles. In this setting, the algorithm is based on factorizing the underlying graph as a modified junction tree and solving an OT problem that is structured according to this modified junction tree. Thus, in this case the problem is posed as a sum of several coupled MOT problems of small size, where each MOT problem corresponds to a clique, i.e., a collection of nodes, in the modified junction tree.
To summarize, our contributions are the following.
(i) For MOT problems associated with tree structures, we reformulate them as coupled bi-marginal OT problems and introduce a novel algorithm for solving them with local entropy regularization in Sec. 4.1. In our approach, we represent the tree as a bipartite graph and categorize the optimization variables into two groups. This partitioning facilitates the parallelization of our algorithm. This advancement bridges the existing gap in the literature by providing a dedicated algorithmic solution for coupled small OT problems.
(ii) In Sec. 4.2, we conduct a thorough complexity analysis for our algorithm addressing coupled bi-marginal OT problems within tree structures. Contrasting with standard MOT problems that iterate over nodes sequentially, our parallelizable algorithm requires only $\widetilde{\mathcal{O}}\left({|E|^{2}}/{\delta^{2}}\right)$ iterations. Here, $|E|$ denotes the total number of edges, and $\delta$ specifies the desired accuracy. This represents a significant efficiency gain, reducing the number of iterations by a factor of $|E|$ compared to the standard MOT problem approach analyzed in Fan et al. (2022a).
(iii) In Sec. 5, we introduce the notion of a modified junction tree, allowing us to extend the scope of our algorithm from tree structures to general graphs. This extension involves formulating the problem as a sum of coupled small MOT problems and constructing a bipartite graph to enable a parallel algorithm. To validate the efficacy and robustness of our algorithm, we have conducted several numerical experiments, the results of which affirm the algorithm’s applicability to graph structures.
(iv) In Sec. 3, we present a concise framework for MOT problems with global entropy regularization. Extending our previous work Fan et al. (2022a), this framework allows for marginal constraints on the joint distributions. Furthermore, it incorporates the capacity for cost functions to depend on multiple marginals, a feature previously implied but now explicitly defined and integrated into our algorithm.
2 Preliminaries and problem formulation
In this section, we review some background material and introduce the problem we consider in this work. In particular, we first define optimal transport and multi-marginal optimal transport and recap the iterative scaling algorithm for approximately solving these problems. We then introduce MOT problems with graphical structure and provide several motivating examples for considering these types of problems.
2.1 Optimal transport and iterative scaling
2.1.1 Optimal transport
The OT problem seeks an optimal transport plan that moves mass from a source distribution to a target one with minimum cost. In this work, we focus on the discrete OT problem (see, e.g., Villani (2003) for the continuous formulations). Thus, consider two non-negative vectors $\mu_{1}\in{\mathbb{R}}_{+}^{d}$ and $\mu_{2}\in{\mathbb{R}}_{+}^{d}$ with mass $1$, i.e., $\sum_{x_{1}}\mu_{1}(x_{1})=\sum_{x_{2}}\mu_{2}(x_{2})=1$, that describe discrete probability distributions. For simplicity of notation, we assume the marginals have equal dimension $d$ throughout, though the extension to the setting with unequal dimensions is straightforward. Here $\mu_{1}(x)$ and $\mu_{2}(x)$ denote the amount of mass in the source distribution and target distribution, respectively, at location $x$. We also define a cost matrix $C\in{\mathbb{R}}^{d\times d}$, where $C(x_{1},x_{2})$ describes the transport cost of moving a unit mass from point $x_{1}$ to $x_{2}$. Similarly, we describe the transport by a non-negative matrix $B\in{\mathbb{R}}_{+}^{d\times d}$, where the element $B(x_{1},x_{2})$ describes the amount of mass transported from point $x_{1}$ to $x_{2}$. This defines the optimal transport problem
$\min_{B\in{\mathbb{R}}_{+}^{d\times d}}{}$ $\operatorname{trace}(C^{\top}B)\qquad\text{ subject to }$ $B{\bf 1}=\mu_{1},\quad B^{\top}{\bf 1}=\mu_{2},$ (1)
where ${\bf 1}$ denotes a vector of ones of proper dimension. The constraints enforce that $B$ is a feasible transport plan between $\mu_{1}$ and $\mu_{2}$, and the objective function $\operatorname{trace}(C^{\top}B)=\sum_{x_{1},x_{2}}C(x_{1},x_{2})B(x_{1},x_{2})$ describes the total cost of transportation for the given transport plan.
2.1.2 Multi-marginal optimal transport
The MOT problem extends the OT problem 1 to finding a transport plan between multiple distributions $\mu_{1},\dots,\mu_{J}$, where $J>2$. Let $\mu_{j}\in{\mathbb{R}}^{d}$ for $j=1,\dots,J$. Then the multi-marginal transport cost and transport plan are described by $J$-mode tensors ${\bf C}\in{\mathbb{R}}^{d\times\dots\times d}$ and ${\bf M}\in{\mathbb{R}}_{+}^{d\times\dots\times d}$. Here, ${\bf C}(x_{1},x_{2},\ldots,x_{J})$ and ${\bf M}(x_{1},x_{2},\ldots,x_{J})$ denote the cost and amount of transport associated with the locations $x_{1},x_{2},\ldots,x_{J}$. Thus, analogously to 1, the MOT problem is defined as
$\min_{{\mathbf{B}}\in{\mathbb{R}}_{+}^{d\times\dots\times d}}\langle{\bf C},{\mathbf{B}}\rangle\qquad\text{ subject to }P_{j}({\mathbf{B}})=\mu_{j},\text{ for }j\in\Gamma,$ (2)
where $\Gamma\subset\{1,2,\dots,J\}$ is an index set specifying which marginal distributions are given, and the projection on the $j$-th marginal of ${\mathbf{B}}$ is computed as
$P_{j}({\mathbf{B}})=\sum_{x_{1},\dots,x_{j-1},x_{j+1},x_{J}}{\mathbf{B}}(x_{1},\dots,x_{j-1},x_{j},x_{j+1},\dots,x_{J}).$ (3)
In the original MOT formulation (Gangbo and Świech 1998, Pass 2015), constraints are given on all marginal distributions, that is the index set in the constraints is $\Gamma=\{1,2,\dots,J\}$. However, as we will see in the next section, in many applications not all marginal distributions are given, thus in this work $\Gamma$ is typically a strict subset of $\{1,2,\dots,J\}$. Finally, we note that the standard bi-marginal OT problem 1 is a special case of the MOT problem 2 with $J=2$ and $\Gamma=\{1,2\}$.
2.1.3 Iterative scaling algorithm
We note that MOT is a standard linear program. However, as it has $d^{J}$ variables, solving it directly scales exponentially with the number of marginals $J$. The computational burden of solving 2 can be partly alleviated by regularizing the objective with an entropy term. This approach has first been proposed for the bi-marginal problem 1 in Cuturi (2013), Chen et al. (2021b), and has later been adapted to multi-marginal problems (Benamou et al. 2015, Haasler et al. 2021b, c). The entropy regularized MOT problem is the strictly convex optimization problem
$\min_{{\mathbf{B}}\in{\mathbb{R}}^{d\times\dots\times d}}\langle{\bf C},{\mathbf{B}}\rangle+\epsilon{\mathcal{H}}({\mathbf{B}}|{\bf M})\qquad\text{ subject to }P_{j}({\mathbf{B}})=\mu_{j},\text{ for }j\in\Gamma,$ (4)
where $\epsilon>0$ is a small regularization parameter and
${\mathcal{H}}({\mathbf{B}}|{\bf M})=\sum_{x_{1},\dots,x_{J}}{\mathbf{B}}(x_{1},\dots,x_{J})\left(\log\left({\mathbf{B}}(x_{1},\dots,x_{J})\right)-\log\left({\bf M}(x_{1},\dots,x_{J})\right)-1\right).$ (5)
with ${\bf M}(x_{1},\dots,x_{J})=\prod_{j\in\Gamma}\mu_{j}(x_{j})$. The entropy regularized problem 4 can be solved by an iterative scaling algorithm, presented in Algorithm 1. This method is also known as Sinkhorn iterations, especially in the bi-marginal setting 1.
Compute ${\bf K}=\exp(-{\bf C}/\epsilon)$, initialize $u_{1},u_{2},\ldots,u_{J}$ to $\mathbf{1}$
while not converged do
       for $j\in\Gamma$ do
             Update ${\bf U}\leftarrow u_{1}\otimes u_{2}\otimes\dots\otimes u_{J}$
            Update $u_{j}\leftarrow u_{j}\odot\mu_{j}./P_{j}({\bf K}\odot{\bf U}\odot{\bf M})$
return ${\mathbf{B}}={\bf K}\odot{\bf U}\odot{\bf M}$
Algorithm 1 Iterative Scaling Algorithm for MOT
In the following, we briefly sketch how to derive Algorithm 1 as a block coordinate ascent method in the dual of 4, for details we refer the reader to Elvander et al. (2020), Haasler et al. (2021b, c). Namely, utilizing Lagrangian duality theory, one can show that the optimal solution to 4 is of the form
${\mathbf{B}}={\bf K}\odot{\bf U}\odot{\bf M},$ (6)
where $\odot$ denotes element-wise multiplication and the tensors are given by
${\bf K}=\exp(-{\bf C}/\epsilon)\qquad{\bf U}=u_{1}\otimes u_{2}\otimes\dots\otimes u_{J},$ (7)
where the vectors $u_{j}\in\mathbb{R}^{d}$ are given by
$u_{j}=\begin{cases}\exp\left(\frac{\lambda_{j}}{\epsilon}\right),&\text{ if }j\in\Gamma\\ {\bf 1},&\text{ otherwise,}\end{cases}$ (8)
and $\lambda_{j}\in\mathbb{R}^{d}$ is the dual variable corresponding to the constraint $P_{j}({\mathbf{B}})=\mu_{j}$ on the $j$-th marginal. The corresponding dual problem of 4 is
$\max_{\{\lambda_{j},j\in\Gamma\}}-\epsilon\langle{\bf K},{\bf U}\rangle{+\sum_{j\in\Gamma}\lambda_{j}^{\top}\mu_{j},}$ (9)
where ${\bf U}$ depends on the dual variables as in 7-8. Algorithm 1 is then derived as a block coordinates ascent in this dual and is therefore guaranteed to converge globally.
2.2 Graph-structured optimal transport
In this work we consider multi-marginal optimal transport problems with a cost that is composed as a sum of cost terms, which each depend on a subset of the marginals. That is, we consider cost tensors of the form
${\bf C}({\bf x})=\sum_{\alpha\in F}C_{\alpha}({\bf x}_{\alpha}),$ (10)
where $F$ is a set of subsets of the tensors indices, i.e., $\alpha\subset\{1,\dots,J\}$ for each element $\alpha\in F$. If we associate the index set of the transport tensor with a set of vertices, $V=\{1,\dots,J\}$, then the set $F$ can be understood as the set of factors in a factor graph (Haasler et al. 2021c, Fan et al. 2022a). In particular, all subsets $\alpha\in F$ that contain two elements can be illustrated as edges in a graph. Moreover, we allow for constraints that act on projections onto several marginals of the transport tensor ${\mathbf{B}}$. For a set $\alpha\subset V$, we define the projection $P_{\alpha}({\mathbf{B}})=\sum_{x_{j},\ j\in V\backslash\alpha}{\mathbf{B}}(x_{1},\dots,x_{J}).$ Let $F_{\Gamma}\subset F$ denote the set of factors which are constrained, and for $\alpha\in F_{\Gamma}$, let $\boldsymbol{\mu}_{\alpha}\in{\mathbb{R}}_{+}^{d^{|\alpha|}}$ be the given tensor, where the number of modes equals the cardinality $|\alpha|$ of $\alpha$. This let us define the graph-structured MOT problem as
$\min_{{\mathbf{B}}\in{\mathbb{R}}_{+}^{d\times\dots\times d}}\langle{\bf C},{\mathbf{B}}\rangle\qquad\text{ subject to }P_{\alpha}({\mathbf{B}})=\boldsymbol{\mu}_{\alpha},\text{ for }\alpha\in F_{\Gamma},$ (11)
where the cost tensor decouples as in 10, and $F_{\Gamma}\subset F$. Note that the graph structured MOT problem 11 is a special case of the standard MOT problem 2 in terms of the cost function. On the other hand, problem 11 allows for a more general class of constraints than the original problem 2.
Remark 1.
We assume that $F_{\Gamma}\subset F$ for convenience of notation. However, in many cases there is not a cost associated with every factor in $F_{\Gamma}$. In this case we can simply define $C_{\alpha}({\bf x}_{\alpha})=0$ for the corresponding $\alpha\in F_{\Gamma}$. In the examples and applications in this paper, we do not explicitly define these vanishing parts of the cost tensor.
The most widely studied graph-structured MOT problem is the Wasserstein barycenter problem (Agueh and Carlier 2011). However, in a recent line of research, graph-structured MOT problems have been utilized in a wide range of applications. For instance, in Haasler et al. (2021b, a) different control and estimation problems for large systems of indistinguishable agents are modeled as graph-structured MOT problems. Even different classes of agents can be modeled with this framework, e.g., in network flow problems (Haasler et al. 2024) and mean field games (Ringh et al. 2021b, 2023). In the following we briefly review a few more exampled of optimal transport problems that fall into the category of graph-structured MOT. In particular, we note that to the best of our knowledge, the Euler problem and Wasserstein splines have not yet been classified as such. Throughout this paper, we illustrate constrained marginals as gray nodes, and white nodes are estimated in the respective problems.
2.2.1 Generalized Euler flow
One of the first applications of multi-marginal optimal transport, the generalized Euler flow (Benamou et al. 2015, Section 4.3), falls in fact into the category of a graph-structured problem. The Euler flow studies a relaxation of the incompressible Euler equation, where one seeks trajectories of particles between constrained initial and final data, that minimize the kinetic energy (Brenier 1989). Discretizing the problem into $J$ time points, we describe the flow by a $J$-mode multi-marginal optimal transport plan ${\mathbf{B}}\in{\mathbb{R}}_{+}^{d^{J}}$, where each marginal describes the density of the flow at time instance $j$, discretized into $d$ points. Note that the density of an incompressible flow, discretized uniformly over a given space, is described by a uniform distribution. Thus, we constrain each marginal of the transport plan by $P_{j}({\mathbf{B}})=\mu_{j}=\frac{1}{d}\boldsymbol{1}$. In the Euler flow, it is enforced that each particle at time 00 is transported to a given location at time $T$, given by a permutation $\sigma:\{1,\dots,d\}\to\{1,\dots,d\}$. This is modeled through a bi-marginal constraint $P_{1,J}({\mathbf{B}})=\boldsymbol{\mu}_{\{1,J\}}=\Pi_{\sigma}$, where $\Pi_{\sigma}\in{\mathbb{R}}^{d\times d}$ is the permutation matrix for the permutation $\sigma$. The Euler flow problem is thus formulated as the MOT problem
$\min_{{\mathbf{B}}\in{\mathbb{R}}_{+}^{d\times\dots\times d}}$ $\langle{\bf C},{\mathbf{B}}\rangle$ (12)
subject to $P_{j}({\mathbf{B}})=\mu_{j}\text{ ~{}~{}for }j=1,\dots,J;\quad P_{1,J}({\mathbf{B}})=\boldsymbol{\mu}_{\{1,J\}},$
where
${\bf C}(x_{1},\dots,x_{J})=\sum_{j=1}^{J-1}C(x_{j},x_{j+1}),$ (13)
and the elements of the matrix $C\in{\mathbb{R}}^{d\times d}$ contain the squared Euclidean distances between the discretization points, namely, $C_{x_{j},x_{j+1}}=\|x_{j}-x_{j+1}\|^{2}$.
Benamou et al. (2015) also propose an alternative relaxed formulation, where the bi-marginal constraint in 12 is replaced by an additional cost term in 13, and the problem becomes
$\min_{{\mathbf{B}}\in{\mathbb{R}}_{+}^{d\times\dots\times d}}\langle{\bf C},{\mathbf{B}}\rangle\qquad\text{ subject to }P_{j}({\mathbf{B}})=\mu_{j}\text{ for }j=1,\dots,J,$ (14)
where
${\bf C}(x_{1},\dots,x_{J})=\sum_{j=1}^{J-1}C(x_{j},x_{j+1})+C^{\sigma}(x_{1},x_{J}),$ (15)
and the elements of the matrix $C^{\sigma}\in{\mathbb{R}}^{d\times d}$ are given by $C^{\sigma}(x_{1},x_{J})=\|x_{\sigma(1)}-x_{J}\|^{2}$.
Both problems 12-13 and 14-15 can be described by a cyclic graph as illustrated in Fig. 1. In particular, we note that although the authors in Benamou et al. (2015) present only the computational method for the relaxed formulation 14-15, our proposed framework encompasses both formulations.
Figure 1: Illustration of graphical structures in some applications. Gray nodes correspond to fixed marginals, and white nodes are estimated in the problem. The dotted edge between $x_{1}$ and $x_{J}$ in (a) may represent either a bi-marginal constraint, as in 12-13, or a cost interaction, as in 14-15.
(a) Euler flow.
$x_{1}$$x_{2}$$\!\!x_{J\!-\!1\!}\!\!$$x_{J}$
(b) Wasserstein least squares.
$x_{1}$$x_{2}$$x_{J}$$x_{0}$$\!\!x_{\!J\!+\!1\!}\!\!$
(c) Wasserstein splines
$x_{0}$$x_{1}$$\!\!x_{J\!-\!1\!}\!\!$$x_{J}$$v_{0}$$v_{1}$$\!\!v_{J\!-\!1}\!\!$$v_{J}$
2.2.2 Wasserstein least squares
Consider a least square fitting problem in the Wasserstein-2 space. Given a set of probability measures $\rho_{t_{1}},\rho_{t_{2}},\ldots,\rho_{t_{J}}$ with $0\leq t_{1}<t_{2}<\cdots<t_{J}\leq 1$, we aim to find two distributions $\mu_{0}$ and $\mu_{1}$ such that the set of measures is closest to the displacement interpolation between $\mu_{0}$ and $\mu_{1}$ in terms of Wasserstein-2 distance. More specifically, we seek the solution to
$\min_{\mu_{0},\mu_{1}}\sum_{j=1}^{J}W_{2}(\rho_{t_{j}},\mu_{t_{j}})^{2}$ (16)
where $\mu_{t}$ denotes the displacement interpolation connecting $\mu_{0}$ and $\mu_{1}$. This least square problem can be reformulated as
$\min_{\pi,\mu_{0},\mu_{1}}$ $\int c(x_{0},x_{1},\cdots,x_{J},x_{J+1}){\rm{d}}\pi(x_{0},x_{1},\cdots,x_{J},x_{J+1})$ (17)
$P_{j}(\pi)=\rho_{t_{j}},~{}~{}j=1,2,\cdots,J.\qquad P_{0}(\pi)=\mu_{0},\qquad P_{J+1}(\pi)=\mu_{1}.$
By discretizing the integral in 17 to a summation, we can transform the problem into the form presented in 11. To enforce that the optimality of the coupling between $\mu_{0}$ and $\mu_{1}$, we can add a large penalty on it, leading to the modified cost
$c(x_{0},x_{1},\cdots,x_{J},x_{J+1})=\sum_{j=1}^{J}\|x_{j}-(1-t_{j})x_{0}-t_{i}x_{J+1}\|^{2}+\alpha\|x_{0}-x_{J+1}\|^{2}$ (18)
with $\alpha>0$ being a sufficiently large number.
2.2.3 Wasserstein Splines
[…]
$\min_{\pi}\int c(x_{0},v_{0},\cdots,x_{J},v_{J}){\rm{d}}\pi(x_{0},v_{0},\cdots,x_{J},v_{J})\qquad\text{subject to }~{}~{}P_{x_{j}}(\pi)=\rho_{t_{j}},\quad j=0,1,\ldots,J$ (19)
where $c(x_{0},v_{0},\cdots,x_{J},v_{J})=$
$\sum_{i=0}^{J-1}\frac{1}{t_{j+1}-t_{j}}\cdot\left[12\left\|\frac{x_{j+1}-x_{j}}{t_{j+1}-t_{j}}-v_{j}\right\|^{2}-12\left\langle\frac{x_{j+1}-x_{j}}{t_{j+1}-t_{j}}-v_{j},v_{j+1}-v_{j}\right\rangle+4\|v_{j+1}-v_{j}\|^{2}\right].$ (20)
This problem is associated with the graphical model in Fig. 1, where the constraints are on the position nodes $x_{0},x_{1},\ldots,x_{J}$.
3 Multi-marginal optimal transport with global regulatization and probabilistic graphical models
A line of recent works have drawn a connection between entropy-regularized multi-marginal optimal transport problems 4 with a graph-structured cost 10 and probabilistic graphical models (PGMs) (Haasler et al. 2021c, Fan et al. 2022a, Haasler et al. 2021b, Singh et al. 2022, Ringh et al. 2023, Zhou and Parno 2024). In this section we review and generalize these connections and the computational methods for solving these types of optimal transport problems.
3.1 A PGM formulation for multi-marginal optimal transport
A PGM provides a compact representation for the joint distribution of a collection of random vectors, which have dependencies that can be described by a graph (Koller and Friedman 2009). Define the tensors $K_{\alpha}({\bf x}_{\alpha})=\exp(-C_{\alpha}({\bf x}_{\alpha})/\epsilon)$ for $\alpha\in F$, where $C_{\alpha}$ are the cost tensors in 10, and $\epsilon$ is the regularization parameter in 4. Then, ${\bf K}$ in 7 has the form ${\bf K}({\bf x})=\prod_{\alpha\in F}K_{\alpha}({\bf x}_{\alpha})$. Moreover, to define the entropy regularization 5 for the graph-structured problem 11 we define the tensor ${\bf M}$ as ${\bf M}({\bf x})=\prod_{\alpha\in F_{\Gamma}}\boldsymbol{\mu}_{\alpha}({\bf x}_{\alpha})$. The factorizations of the tensors ${\bf K}$ and ${\bf M}$ can be used to define the joint probability distribution of a PGM,
$p({\bf x})=\frac{1}{Z}\left(\prod_{\alpha\in F}K_{\alpha}({\bf x}_{\alpha})\right)\left(\prod_{\alpha\in F_{\Gamma}}\boldsymbol{\mu}_{\alpha}({\bf x}_{\alpha})\right),$ (21)
where $Z$ is a normalization constant. Moreover, note that the solution tensor 6 can be written as ${\bf K}\odot{\bf U}\odot{\bf M}$ with ${\bf U}({\bf x})=\prod_{\alpha\in F_{\Gamma}}{\mathbf{u}}_{\alpha}({\bf x}_{\alpha})$. Thus, the scaling factors ${\mathbf{u}}_{\alpha}$, for $\alpha\in F_{\Gamma}$ can be interpreted as local potentials of the modified graphical model ${\bf K}({\bf x}){\bf U}({\bf x}){\bf M}({\bf x})$ The Iterative Scaling algorithm 1 finds potentials ${\mathbf{u}}_{\alpha}$, for $\alpha\in F_{\Gamma}$, such that the graphical model defined by the tensor ${\mathbf{B}}={\bf K}\odot{\bf U}\odot{\bf M}$ satisfies all the constraints $P_{\alpha}({\mathbf{B}})=\boldsymbol{\mu}_{\alpha}$, for $\alpha\in F_{\Gamma}$.
3.2 Iterative Scaling Belief Propagation
Each step in the Iterative Scaling algorithm involves calculating a projection $P_{j}({\bf K}\odot{\bf U}\odot{\bf M})$, which in terms of the PGM means inferring the distribution of the $j$-th variable. This Bayesian inference problem can be addressed by the Belief Propagation algorithm (Mezard and Montanari 2009, Part D), (Koller and Friedman 2009, Sec. 20.5.1). Belief propagation is known to converge when the underlying graph is a tree, i.e., does not have any loops, and reduces the complexity of computing the projections in the Iterative Scaling algorithm immensely. More precisely, whereas a brute force approach to compute the projections 3 scales exponentially, Belief Propagation scales only linearly in the number of marginals (Haasler et al. 2021c).
In Fan et al. (2022a), a similar approach has be utilized for general graphical models with possible loops, by partitioning the graph as a so-called junction tree. A junction tree (also called tree decomposition) describes a partitioning of a graph, where several nodes are clustered together, such that the interactions between the clusters can be described by a tree structure.
Definition 1 (Junction tree (Dawid 1992)).
A junction tree $\mathcal{T}=(\mathcal{C},\mathcal{E})$ over a graph $G=(V,F)$ is a tree whose nodes are cliques $c\in\mathcal{C}$ associated with subsets ${\bf x}_{c}\subset V$, and satisfying the properties:
  • Family preservation: For each $\alpha\in F$ there is a clique $c\in{\mathcal{C}}$ such that $\alpha\subset c$.
  • Running intersection: For every pair of cliques $c_{i},c_{j}\in\mathcal{C}$, every clique on the path between $c_{i}$ and $c_{j}$ contains $c_{i}\cap c_{j}$.
For two adjoining cliques $c_{i}$ and $c_{j}$, define the separation set $s_{ij}=\{v\in V:v\in c_{i}\cap c_{j}\}$.
A graph typically has several possible junction trees. In particular, one trivial junction tree is always given by a single cluster containing all nodes. For our computational method, we set up the junction tree such that each constraint marginal or clique is represented by a leaf in the junction tree. The entropy-regularized graph structured optimal transport problem with junction tree decomposition $\mathcal{T}=(\mathcal{C},\mathcal{E})$ is thus of the form
$\min_{{\mathbf{B}}\in{\mathbb{R}}_{+}^{d\times\dots\times d}}\langle{\bf C},{\mathbf{B}}\rangle+\epsilon{\mathcal{H}}({\mathbf{B}}{|{\bf M}})\qquad\text{ subject to }P_{c}({\mathbf{B}})=\boldsymbol{\mu}_{c},\text{ for }c\in\mathcal{C}_{\Gamma},$ (22)
where $\mathcal{C}_{\Gamma}\subset\mathcal{C}$ is the set of constrained cliques, and the cost tensor decouples as
${\bf C}({\bf x})=\sum_{c\in\mathcal{C}}{\bf C}_{c}({\bf x}_{c}).$ (23)
Such a junction tree representation exists for any graph, however it may require adding artificial clusters to the junction tree decomposition.
Example 1.
Some examples of valid junction trees for our method are shown in Fig. 2 (for the Euler flow example in Fig. 1), Fig. 3 (for the Wasserstein least squares problem 17 in Fig. 1) , and Fig. 4 (for the Wasserstein splines problem in Fig. 1).
Figure 2: Junction tree for the graph from the Euler flow problem in Fig. 1, with leaf nodes corresponding to the constrained marginals.
[…]
Figure 3: Junction tree for the graph for the Wasserstein least square problem from Fig. 1. The junction tree is constructed such that the leaves of the tree correspond to the constraints in the least square problem 17.
[…]
Figure 4: Junction tree of Wasserstein splines in Fig. 1.
$\begin{subarray}{c}v_{0},v_{1}\\ x_{0},x_{1}\end{subarray}$$\begin{subarray}{c}v_{1}\\ x_{1}\end{subarray}$$\begin{subarray}{c}v_{1},v_{2}\\ x_{1},x_{2}\end{subarray}$$\begin{subarray}{c}v_{J-2}\\ v_{J-1}\\ x_{J-2}\\ x_{J-1}\end{subarray}$$\begin{subarray}{c}v_{J-1}\\ x_{J-1}\end{subarray}$$\begin{subarray}{c}v_{J-1},v_{J}\\ x_{J-1},x_{J}\end{subarray}$$x_{0}$$x_{1}$$x_{J\!-\!2\!}$$x_{J\!-\!1\!},x_{J}$$x_{0}$$x_{1}$$x_{J\!-\!2}$$\begin{subarray}{c}x_{J-1}\\ x_{J}\end{subarray}$
Input: Given a junction tree $\mathcal{T}=(\mathcal{C},\mathcal{E})$ with constraints on the leaf cliques $\mathcal{C}_{\Gamma}\subset\mathcal{C}$
Initialize messages $m_{i\to j}(s_{ij}),\ m_{j\to i}(s_{ji}),\ \forall(i,j)\in\mathcal{E}$
while not converged do
$m_{i\to j}(s_{ij})$ $\leftarrow\sum_{x_{c_{i}}\setminus s_{ij}}{\bf K}_{c_{i}}({\bf x}_{c_{i}})\prod_{\ell\in N(i)\setminus j}m_{\ell\to i}(s_{\ell i}),\qquad c_{i}\in{\mathcal{C}}\setminus{\mathcal{C}}_{\Gamma}$ (24)
$m_{i\to j}(s_{ij})$ $\leftarrow\mu_{c_{i}}({\bf x}_{c_{i}})/m_{j\to i}({\bf x}_{c_{i}}),\qquad c_{i}\in{\mathcal{C}}_{\Gamma}.$ (25)
return $m_{i\to j}(s_{ij}),\ m_{j\to i}(s_{ji}),\ \forall(i,j)\in\mathcal{E}$
Algorithm 2 Iterative Scaling Belief Propagation Algorithm for graphical MOT with global regularization
Similarly to the factorization 21, we can then write the probability distribution over the PGM following the junction tree decomposition $\mathcal{T}=(\mathcal{C},\mathcal{E})$ as a product
$p({\bf x})=\frac{1}{Z}\left(\prod_{c\in\mathcal{C}}{\bf K}_{c}({\bf x}_{c})\right)\left(\prod_{c\in\mathcal{C}_{\Gamma}}\boldsymbol{\mu}_{c}({\bf x}_{c})\right).$ (26)
Let $N(i)$ be the neighboring nodes of the node $i$. By extending the Iterative Scaling Belief Propagation methods in Haasler et al. (2021c), Fan et al. (2022a), we propose Algorithm 2 to solve 22. Upon convergence of Algorithm 2, the optimal transport plans on the clusters are given by
${\mathbf{B}}_{c}({\bf x}_{c})={\bf K}_{c}({\bf x}_{c})\prod_{\ell\in N(i)}m_{\ell\to i}(s_{\ell i}).$ (27)
The algorithm sends messages between neighboring cliques. The messages are thus of the same size as the separation sets. Recall that the leaves of the junction tree are associated with constraints in the optimal transport problem. Thus, the messages sent from these leaves correspond to an update step similar to the one in Algorithm 1. The messages sent within the graph encode the computation of the projections. More precisely, picking one leaf $c$ and sending the messages from all other leaves across the separation sets towards our leaf amounts to one exact computation of the projection $P_{c}({\bf K}\odot{\bf U}\odot{\bf M})$. However, the algorithm converges with any message schedule (Haasler et al. 2021c, Fan et al. 2022a).
Note that in the case where all cost are pair-wise and constraints act on one-marginals of the transport tensors, Algorithm 2 boils down to the Iterative Scaling Belief Propagation algorithms in Haasler et al. (2021c), Fan et al. (2022a). Here, we have extend on these methods and provide a more general algorithm for the case where we may have constraints on joint distributions, and where cost terms can depend on more than two marginals.
The Iterative Scaling Belief Propagation algorithm utilizes the graphical structure in problem 22-23 to significantly improve on the complexity of the Sinkhorn iterations. Recall that the computational bottleneck of Algorithm 1 lies in the computation of the projections, which require $\mathcal{O}(d^{J})$ operations. Algorithm 2 reduces this complexity by utilizing the graphical structure for computing the projections. The complexity of the problem depends then on the size of the messages, i.e., the size of the separation sets. This quantity can also be described by the so-called tree-width.
Definition 2 (Tree width (Bertele and Brioschi 1973, Halin 1976)).
For a junction tree $\mathcal{T}=(\mathcal{C},\mathcal{E})$, we define its width as $\text{width}({\mathcal{T}})=\max_{c\in{\mathcal{C}}}|c|-1$. For a graph $G$, we define its tree-width as $w(G)=\min\{\text{width}({\mathcal{T}})\ |\ {\mathcal{T}}\text{ is a junction tree for }G\}$.
In order for Algorithm 2 to be efficient we thus would like to pick a junction tree factorization with small tree-width. Given a graph, a junction tree decomposition that minimizes the tree-width is also called a minimal junction tree. In general, finding a junction tree with minimal width is a NP hard problem (Koller and Friedman 2009). However, for all applications in this paper it is easy to find minimal junction trees.
Example 2.
The junction trees in Figures 2, 3, and 4 are minimal junction trees for the respective graphs. Since the cliques in the former two junction trees are of cardinality 3, the respective graphs’ tree-widths are 2. Similarly, the last graph’s tree width is 3.
For Algorithm 2, we can directly apply the complexity analysis in Fan et al. (2022a). We will discuss the complexity result in remark 2.
4 Tree-structured multi-marginal optimal transport with local regularization
In the context where the MOT problem is tree structured, i.e., structured as a graph without cycles, the innate characteristics of tensor inner products facilitate a natural reformulation of the problem. This reformulation manifests as a sum of coupled bi-marginal OT problems. Considering a set of fixed marginals, denoted as $\mu_{j}$ for each $j$ belonging to the vertex set $\Gamma$, the problem of coupled bi-marginal OT can be written as:
$\min_{\begin{subarray}{c}B_{(j,k)},(j,k)\in E\\ \mu_{j},j\in V\setminus\Gamma\end{subarray}}\sum_{(j,k)\in E}{\langle}C_{(j,k)},B_{(j,k)}{\rangle}\qquad\text{subject to}\quad B_{(j,k)}\mathbf{1}=\mu_{j},~{}~{}B_{(j,k)}^{\top}\mathbf{1}=\mu_{k}~{}~{}\text{for }(j,k)\in E,$ (28)
where $E$ is a set of all the edges of the graph. The matrices $C_{(j,k)}\in\mathbb{R}^{d\times d}$ and $B_{(j,k)}\in\mathbb{R}^{d\times d}$ represent the transport cost and transport plan corresponding to the edge connecting nodes $j$ and $k$, respectively. We implicitly assume the convention $B_{(j,k)}=B_{(k,j)}^{\top}$ and $C_{(j,k)}=C_{(k,j)}^{\top}$. In contrast to the global entropy regularization presented in Sec. 3 for a standard MOT problem 22, applying local regularization on each edge appears as a more direct and natural approach within this framework. Notably, in tree-structured optimal transport scenarios, such as barycenter problems, this form of regularization is frequently employed, as described in existing literature (Kroshnin et al. 2019).
4.1 Algorithm
Drawing inspiration from Kroshnin et al. (2019), we developed a parallel algorithm to address the coupled bi-marginal OT problem for general trees. Herein, we propose to use a modified entropy regularization similar to Fan et al. (2022a). The reference joint distribution between nodes $j$ and $k$ is denoted as:
$M_{(j,k)}:=\gamma_{j}\otimes\gamma_{k},$ (29)
where $\gamma_{j}=\mu_{j}$ if $j\in\Gamma$ and $\mathbf{1}$ otherwise. As discussed before, we always assume that marginals may only be constrained if they are on the leaves of the tree. It is important to note that therefore the scenario where $(j,k)\in E$ and both $j,k\in\Gamma$ does not exist. In such cases, the OT problem simplifies into several independent subproblems. Consequently, this leads us to formulate the regularized problem with an enforced mass constraint.
$\min_{\begin{subarray}{c}B_{(j,k)},(j,k)\in E\\ \mu_{j},j\in V\setminus\Gamma\end{subarray}}$ $\sum_{(j,k)\in E}\langle C_{(j,k)},B_{(j,k)}\rangle+\epsilon{\mathcal{H}}(B_{(j,k)}|M_{(j,k)})$ (30)
subject to $B_{(j,k)}\mathbf{1}=\mu_{j},~{}B_{(j,k)}^{\top}\mathbf{1}=\mu_{k}\text{ for }(j,k)\in E;~{}~{}~{}\mathbf{1}^{\top}\mu_{j}=1\text{ for }j\in V\backslash\Gamma.$
The regularization, as formulated, is applied locally to each edge of the tree, in contrast to the previous formulations 22, which regularizes the multi-marginal optimal transport tensor. Utilizing a standard Lagrangian dual theory enables the derivation of the dual problem. Denote
$B_{(j,k)}=\operatorname{diag}(u_{(j,k)}\odot\gamma_{j})K_{(j,k)}\operatorname{diag}(u_{(k,j)}\odot\gamma_{k}),\qquad K_{(j,k)}=\exp(-C_{(j,k)}/\epsilon)$ (31)
where $u_{(j,k)}\in{\mathbb{R}}^{d}$ Furthermore, $u=\{u_{(j,k)}\}_{j\neq k}$ and $\rho=\{\rho_{j}\}_{j\in V\backslash\Gamma}$. For a node $j\in\Gamma$, its unique neighbor is denoted by $k_{j}$. The dual problem is then formulated as:
$\min_{u,\rho}~{}~{}f(u,\rho):=$ $\sum_{(j,k)\in E}\|B_{(j,k)}\|_{1}-\sum_{j\in\Gamma}\log(u_{(j,k_{j})})^{\top}\mu_{j}-\sum_{j\in V\backslash\Gamma}\rho_{j}$ (32)
subject to $\ \sum_{k\in N(j)}\log(u_{(j,k)})=\rho_{j}\mathbf{1},\quad\text{for }j\in V\backslash\Gamma.$
We then introduce another set of dual variables and apply the Lagrangian dual theory again to the dual problem 32, and obtain the update rules. The dual variables we introduce are reflected in $v_{j}$ in 34.
Theorem 1.
The optimal dual variables can be found as the limit point of the following iterative scheme: For $j\in\Gamma$ we update
$u_{(j,k_{j})}=\mathbf{1}./(K_{(j,k_{j})}u_{(k_{j},j)}).$ (33)
For $j\in V\setminus\Gamma$ we update the set of vectors $u_{(j,k)}$, where $k\in N(j)$, according to
$u_{(j,k)}=e^{\rho_{j}/|N(j)|}v_{j}./(K_{(j,k)}(u_{(k,j)}\odot\gamma_{k})),\qquad v_{j}=\bigg{(}\bigodot_{k\in N(j)}\Big{(}K_{(j,k)}(u_{(k,j)}\odot\gamma_{k})\Big{)}\bigg{)}^{1/|N(j)|}.$ (34)
and we update $\rho_{j}=-|N(j)|\log\left(\mathbf{1}^{\top}v_{j}\right).$
Based on Theorem 1, we note that for any node $j$, the updates of $\{u_{(j,k)}\}_{k\in N(j)}$ or $\rho_{j}$ depend only on the variables $u_{(k,j)}$ from neighboring nodes $k\in N(j)$. This observation is critical for enabling parallel updates in the algorithm, as it allows for independently updating dual variables associated with two distinct parts of the tree. The concept of a bipartite graph becomes instrumental here, providing the framework for such parallelism.
Definition 3.
(Bipartite Graph (Metcalf and Casey 2016)). A bipartite graph is a graph where the vertices can be divided into two disjoint sets such that all edges connect a vertex in one set to a vertex in another set. There are no edges between vertices in the disjoint sets.
Figure 5: Bipartite graph with sets $S_{1},$ $S_{2}$.
[…]
To leverage this, the algorithm unfolds in the following manner: We begin by representing the tree as a bipartite graph, a natural representation since all trees are inherently bipartite according to Scheinerman (2012). The algorithm then iteratively updates until convergence, alternating between: i) Updating all dual variables from the first partition to the second. ii) Updating all dual variables from the second partition to the first. Denote the intermediate computational variables
$q^{t}_{(j,k)}=B^{t}_{(j,k)}\mathbf{1},\qquad q^{t+1}_{j}=\bigg{(}\bigodot_{k\in N(j)}q^{t}_{(j,k)}\bigg{)}^{\frac{1}{|N(j)|}}.$ (35)
Our method is detailed in Algorithm 3, termed Bipartite Iterative Scaling. A supporting proposition 1 underpins its theoretical foundation.
Input: Probability vectors $\{\mu_{j}\}_{j\in\Gamma}$. Stopping criteria $\delta^{\prime}$. Transport cost $\{C_{(j,k)}\}$. Regularization param. $\epsilon$.
Construct a bipartite graph where the nodes are divided into two disjoint groups $\{S_{1},S_{2}\}$.
Initialization of variables: $t=1,~{}~{}u^{0}_{(j,k)}=\mathbf{1}\in{\mathbb{R}}^{d}$ for $j\neq k$.
while $\sum_{j\in\Gamma}\|B_{(j,k_{j})}^{t}\mathbf{1}-\mu_{j}\|_{1}+\sum_{j\notin\Gamma}\sum_{k\in N(j)}\|q_{(j,k)}^{t}-\bar{q}_{j}^{t}\|_{1}\geq\delta^{\prime}$  do
       Let $S=S_{1}$ if $t$ is odd, otherwise $S=S_{2}$
      Calculate $q_{\{j,k\}}^{t},~{}~{}q_{j}^{t+1}$ according to 35 for $j\in S$
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!u^{t+1}_{(j,k_{j})}=u^{t}_{(j,k_{j})}\odot\mu_{j}./q^{t}_{(j,k_{j})}\quad\text{for }j\in\Gamma\bigcap S$ (36)
$\!\!\!\!\!\!\!\!\!\!\!\!u_{(j,k)}^{t+1}=u_{(j,k)}^{t}\odot\frac{q_{j}^{t+1}}{\|q_{j}^{t+1}\|_{1}}./q_{(j,k)}^{t}\quad\text{for }j\in(V\backslash\Gamma)\bigcap S$ (37)
      $t\leftarrow t+1$
      Construct ${B}^{t}_{(j,k)},~{}(j,k)\in E$ according to 31.
return ${B}^{t}_{(j,k)}$ for all $(j,k)\in E$
Algorithm 3 Bipartite Iterative Scaling Algorithm for graphical coupled bi-marginal OT with local regularization
/* Post-update of nodes in one partition, the projections on nodes in the opposite partition do not match the marginals. This partition is denoted as $S$. */
Input: The transportation plans ${B}_{(j,k)}$ for all $(j,k)\in E$, where $P_{(j,k)}$ has mismatch with marginal $q_{j}$ when $j\in S$. A sequence of probability vectors $\{\mu_{j}\}_{j\in\Gamma}$
for $j\in S$ do
       if $j\in\Gamma$ then
            Round $B_{(j,k_{j})}$ by Algorithm 2 in Altschuler et al. (2017) such that $\widehat{B}_{(j,k_{j})}\in\Pi(\mu_{j},B^{\top}_{(j,k_{j})}{\bf 1})$
      else
            Round $B_{(j,k)}$ by Algorithm 2 in Altschuler et al. (2017) such that $\widehat{B}_{(j,k)}\in\Pi(\bar{q}_{j}(B),B^{\top}_{(j,k)}{\bf 1})$ for $\forall k\in N(j)$, where $\bar{q}_{j}(B)=\frac{1}{|N(j)|}\sum_{k\in N(j)}B_{(j,k)}\mathbf{1}$.
return $\widehat{B}_{(j,k)}$ for all $(j,k)\in E$
Algorithm 4 Rounding for transport plan on a bipartite graph
Proposition 1.
The updates in Theorem 1 can be equivalently expressed as the updates 36-37 in Algorithm 3 together with the variables 35.
The updates of the dual variables, as specified in 36-37, are independent within the same group $S$, facilitating simultaneous updates in parallel. Provided there is sufficient memory to store all variables, concurrent updating is feasible.
In each iteration Algorithm 3 updates the variables corresponding to one of the partitions $S_{1}$ or $S_{2}$. This update guarantees that the constraints are matched exactly for the nodes in this partition. However, this means that after any finite number of iterations of Algorithm 3 the marginal constraints will not be matched for the nodes in the partition that has not been updated in the last iteration. To address this, we adopt a strategy akin to the ones in Altschuler et al. (2017), Fan et al. (2022a) and employ a Rounding Algorithm 4. This ensures the generation of transport plans that adhere to all stipulated constraints in 30. The complexity of the rounding process is ${\mathcal{O}}(|E|d^{2})$. Algorithm 4 differs from the method in Fan et al. (2022a) in that it requires modification of the transport plans across all edges.
4.2 Complexity analysis
In this section, we provide a rigorous demonstration that our Algorithm 3, when used in conjunction with the Rounding Algorithm 4, requires $T=\widetilde{\mathcal{O}}\left({|E|^{2}}/{\delta^{2}}\right)$ iterations to achieve a $\delta$-approximation of the problem 28. First, we will illustrate that, through the specific mass normalization technique employed in 37, the transport plan maintains unit mass in each iteration.
Lemma 1.
At each update $t$ of Algorithm 3, we have $\|B^{t}_{(j,k)}\|_{1}=\|q_{(j,k)}^{t}\|_{1}=1,~{}~{}\forall(j,k)\in E.$
We now define the variable $\lambda^{t}_{(j,k)}=\epsilon\log(u^{t}_{(j,k)})$ for $(j,k)\in E$, and $u^{t}_{(j,k)}$ are the iterates in Algorithm 3. Lemma 2 provides bounds on the scaling variables.
Lemma 2.
Denote $C_{\infty}:=\max_{(j,k)\in E}\|C_{(j,k)}\|_{\infty}.$ For all iterates $\lambda_{(j,k)}^{t}$, and their fixed points $\lambda_{(j,k)}^{*}$,
$\max_{i=1,\dots,d}(\lambda_{(j,k)}^{t})_{i}-\min_{i=1,\dots,d}(\lambda_{(j,k)}^{t})_{i}\leq 2C_{\infty},\qquad\max_{i=1,\dots,d}(\lambda_{(j,k)}^{*})_{i}-\min_{i=1,\dots,d}(\lambda_{(j,k)}^{*})_{i}\leq 2C_{\infty}.$ (38)
The following lemma links the dual objective error in Theorem 1 to the stopping criteria in Algorithm 3.
Lemma 3.
For all iterates $(u^{t},\rho^{t})$ of the method in Algorithm 3 and their fixed points $(u^{*},\rho^{*})$, it holds that
$f(u^{t},\rho^{t})-f(u^{*},\rho^{*})\leq\frac{2C_{\infty}}{\epsilon}\left(\sum_{j\in{\Gamma}}\|B_{(j,k_{j})}^{t}\mathbf{1}-\mu_{j}\|_{1}+\sum_{j\in{V\backslash\Gamma}}\sum_{k\in N(j)}\|B_{(j,k)}^{t}\mathbf{1}-\bar{q}_{j}^{t}\|_{1}\right)$ (39)
where $\bar{q}_{j}^{t}=\frac{1}{|N(j)|}\sum_{k\in N(j)}q_{(j,k)}^{t}=\frac{1}{|N(j)|}\sum_{k\in N(j)}B_{(j,k)}^{t}\mathbf{1}$.
The increase in the dual objective between iterations is also closely related to the stopping criteria in Algorithm 3.
Lemma 4.
For the iterates in Algorithm 3, it holds that
$f(u^{t},\rho^{t})-f(u^{t+1},\rho^{t+1})\geq\frac{1}{22|E|}\left(\sum_{j\in{\Gamma}}\|B_{(j,k_{j})}^{t}\mathbf{1}-\mu_{j}\|_{1}+\sum_{j\in{V\backslash\Gamma}}\sum_{k\in N(j)}\|B_{(j,k)}^{t}\mathbf{1}-\bar{q}_{j}^{t}\|_{1}\right)^{2}.$ (40)
Unlike Kroshnin et al. (2019), our Lemma 3 and Lemma 4 apply universally for any $t$, obviating the need to consider odd and even $t$ distinctly.
Theorem 2.
Algorithm 3 generates $\{B_{(j,k)}^{t}\}$ satisfying $\sum_{j\in\Gamma}\|B_{(j,k_{j})}^{t}\mathbf{1}-\mu_{j}\|_{1}+\sum_{j\notin\Gamma}\sum_{k\in N(j)}\|q_{(j,k)}^{t}-\bar{q}_{j}^{t}\|_{1}\leq\delta^{\prime}$ within $t=2+\frac{88|E|C_{\infty}}{\delta^{\prime}\epsilon}$ iterations.
The subsequent two lemmas quantify the error introduced to the objective value by the rounding process.
Lemma 5.
Let $\{B_{(j,k)}\}$ be a sequence of transportation plans on a bipartite tree. Algorithm 4 returns $\{\widehat{B}_{(j,k)}\}$ satisfying $\widehat{B}_{(j,k)}^{\top}{\bf 1}=q_{j}$ for $j\in V$, where $q_{j}$ is the marginal distribution of node $j$. Moreover, it holds that $\bigg{|}\sum_{(j,k)\in E}{\langle}C_{(j,k)},B_{(j,k)}{\rangle}-\sum_{(j,k)\in E}{\langle}C_{(j,k)},\widehat{B}_{(j,k)}{\rangle}\bigg{|}\leq 2C_{\infty}\left(\sum_{j\in\Gamma}\|\mu_{j}-B_{(j,k_{j})}{\bf 1})\|_{1}+\sum_{j\notin\Gamma}\sum_{k\in N(j)}\|\bar{q}_{j}(B)-B_{(j,k)}{\bf 1}\|_{1}\right),$ where $\bar{q}_{j}(B)=\frac{1}{|N(j)|}\sum_{k\in N(j)}B_{(j,k)}\mathbf{1}$.
Lemma 6.
Let $\{\widetilde{B}_{(j,k)}\}$ be the output of Algorithm 3 and let $\{\widehat{B}_{(j,k)}\}$ be the output of Algorithm 4 with input $\{\widetilde{B}_{(j,k)}\}$, and denote $\{B^{*}_{(j,k)}\}$ the optimal solution to the unregularized MOT problem 28. Then it holds that
$\sum_{(j,k)\in E}{\langle}C_{(j,k)},\widehat{B}_{(j,k)}{\rangle}-\sum_{(j,k)\in E}{\langle}C_{(j,k)},B^{*}_{(j,k)}{\rangle}$ (41a)
$\leq$ $2\epsilon|E|\log(d)+4C_{\infty}\left(\sum_{j\in\Gamma}\|\mu_{j}-\widetilde{B}_{(j,k_{j})}{\bf 1})\|_{1}+\sum_{j\notin\Gamma}\sum_{k\in N(j)}\|\bar{q}_{j}(\widetilde{B})-\widetilde{B}_{(j,k)}{\bf 1}\|_{1}\right).$ (41b)
Given a predetermined final error threshold $\delta$, we select parameters $\epsilon=\frac{\delta}{4|E|\log(d)}$ and $\delta^{\prime}=\frac{\delta}{8C_{\infty}}$ for use in Algorithm 3. This results in an output of $\widetilde{B}_{(j,k)}$, for each edge pair $(j,k)\in E$. Next, we employ Algorithm 4 to round $\widetilde{B}_{(j,k)}$ for each $(j,k)\in E$, thereby producing $\widehat{B}_{(j,k)}$. With these procedures in place, we are equipped to compute the final complexity for the coupled bi-marginal OT problem as described in 28.
Theorem 3.
Algorithm 3, in conjunction with Algorithm 4, yields a $\delta$-approximate solution to the tree-structured MOT problem 28 in $T$ arithmetic operations, where $T=\mathcal{O}\left(\frac{|E|^{3}C_{\infty}^{2}d^{2}\log(d)}{\delta^{2}}\right).$
Remark 2 (Comparison of complexity with global regularization (Fan et al. 2022a)).
We compare our complexity results with the ones in Fan et al. (2022a), for graph-structured MOT with global entropic regularization in terms of the total complexity and the number of iterations. i) Total Complexity: According to Theorem 2 in Fan et al. (2022a), global entropy regularization requires $\widetilde{\mathcal{O}}(d(G)|E||\Gamma|^{2}C_{\infty}^{2}d^{2}/\delta^{2})$ arithmetic operations with high probability, where $d(G)$ denotes the maximum node distance in graph $G$. Under the assumptions $|E|={\mathcal{O}}(|\Gamma|)$ and $d(G)={\mathcal{O}}(1)$, akin to the Wasserstein barycenter problem, both the global and local regularization methods offer similar complexity bounds. However, a notable distinction is that Theorem 3 offers a deterministic bound. ii) Iterations: Algorithm 3 with local regularization, as per Theorem 2 and Theorem 3, necessitates $\mathcal{O}\left(\frac{|E|^{2}C_{\infty}^{2}\log(d)}{\delta^{2}}\right)$ iterations, less than the requirement for global regularization at ${\mathcal{O}}\left(\frac{|E||\Gamma|^{2}C_{\infty}^{2}\log(d)}{\delta^{2}}\right)$. The reduced iteration count is a benefit of algorithm parallelization, albeit with a higher computational load per iteration.
5 Extension of the algorithm with local regularization to general
graphs
Building on the concepts from Sec. 4, we extend our approach to OT problems with a general factor graph structure as indicated in 10. We employ a strategy that breaks down the MOT problem into interconnected smaller MOT problems, introducing local regularization for each sub-problem. To facilitate this, we define a specialized type of junction tree, elaborated in Definition 1.
Definition 4.
A modified junction tree is a class of junction tree $\mathcal{T}=(\mathcal{C},\mathcal{E})$ with two disjoint sets of cliques ${\mathcal{Q}}_{C}\subset{\mathcal{C}}$ and separators ${\mathcal{Q}}_{S}\subset{\mathcal{C}}$ such that
  1. (1)
    For any $c\in{\mathcal{Q}}_{S}$, its neighbors $N(c)\subseteq{\mathcal{Q}}_{C}$, and for any $c\in{\mathcal{Q}}_{C}$, its neighbors $N(c)\subseteq{\mathcal{Q}}_{S}$.
  2. (2)
    Family preservation: For each $\alpha\in F$ there is a clique $c\in{\mathcal{Q}}_{C}$ such that $\alpha\subset{c}$.
  3. (3)
    Running intersection: For every pair of cliques $c_{i},c_{j}\in{\mathcal{Q}}_{C}$, every clique in ${\mathcal{Q}}_{C}$ or ${\mathcal{Q}}_{S}$ on the path between $c_{i}$ and $c_{j}$ contains $c_{i}\cap c_{j}$.
  4. (4)
    One clique is only connected to two separators: For all $c\in{\mathcal{Q}}_{C}$ it holds $|N(c)|=2$.
Note that each separator set in Definition 1 is connected to exactly two cliques. On the other hand, by Definition 4, our cliques in ${\mathcal{Q}}_{S}$ can be connected to an arbitrary number of cliques, and instead, the cost-cliques in ${\mathcal{Q}}_{C}$ can only be connected to two cliques. This latter assumption is necessary for efficient parallelization of the algorithm we introduce and can be relaxed as we will explain in remark 3.
Example 3.
The junction trees in Figures 6, 7, and 8 are modified junction trees for the respective graphs. In contrast to junction trees in Figures 2, 3, and 4, each clique (shown as circles) can only be connected to two separators (shown as squares), while separator can connect to multiple cliques. Moreover, note that constraints in the respective MOT problems are indicated by gray boxes in Figures 6, 7, and 8. This is explained in more detail in the following section.
Figure 6: Modified junction tree for the graph of Euler flow example from Fig. 1.
[…]
Figure 7: Modified junction tree for the graph of Wasserstein least square example from Fig. 1.
[…]
Figure 8: Modified junction tree for the graph of Wasserstein spline example from Fig. 1.
$\begin{subarray}{c}v_{0},v_{1}\\ x_{0},x_{1}\end{subarray}$$x_{1}$$v_{1}$$\begin{subarray}{c}v_{1},v_{2}\\ x_{1},x_{2}\end{subarray}$$x_{2}$$v_{2}$$\begin{subarray}{c}v_{2},v_{3}\\ x_{2},x_{3}\end{subarray}$$\begin{subarray}{c}v_{J-2},v_{J-1}\\ x_{J-2},x_{J-1}\end{subarray}$$x_{J-1}$$v_{J-1}$$\begin{subarray}{c}v_{J-1},v_{J}\\ x_{J-1},x_{J}\end{subarray}$$x_{J}$$x_{0}$
5.1 Problem formulation
We structure a coupled MOT problem 42, which is equivalent to MOT problem 11, on the junction tree in Definition 4 by associating every cost term with a clique in ${\mathcal{Q}}_{C}$. Moreover, we assume that each constraint is embedded in a separator in ${\mathcal{Q}}_{S}$ as described in the following. A constraint of the optimal transport problem can act on projections of the solution tensors ${\mathbf{B}}_{c}$ onto a set of marginals. Recall that $\Gamma\subset\{1,2,\dots,J\}$ denote the set of index-sets corresponding to the constraints (see 2). Then we assume that there exists a partition of $\Gamma$ such that for each ${\boldsymbol{\gamma}}$ in the partition there is a $c_{\boldsymbol{\gamma}}\in{\mathcal{Q}}_{S}$ such that ${\boldsymbol{\gamma}}\subseteq c_{\boldsymbol{\gamma}}$, and any $c_{\boldsymbol{\gamma}}$ can contain only one constrained set ${\boldsymbol{\gamma}}$. Moreover, we denote the separators with constraints as ${\mathcal{Q}}_{\Gamma}=\{c_{\boldsymbol{\gamma}}\}{\subseteq{\mathcal{Q}}_{S}}$. Note that ${\boldsymbol{\gamma}}$ may not only contain one node, for example, if $c_{\boldsymbol{\gamma}}=\{x_{1},x_{2},x_{3}\}$, we allow ${\boldsymbol{\gamma}}=\{x_{1},x_{3}\}$. Thus, we define the coupled MOT problem
$\min_{{\mathbf{B}}_{c},c\in{\mathcal{Q}}}$ $\sum_{c\in{\mathcal{Q}}_{C}}{\langle}{\bf C}_{c},{\mathbf{B}}_{c}{\rangle}$ (42)
subject to $P_{{c_{i}}\cap{c_{j}}}({\mathbf{B}}_{c_{i}})=P_{{c_{i}}\cap{c_{j}}}({\mathbf{B}}_{c_{j}}),\quad\text{ for each }c_{i}\in{\mathcal{Q}}_{S},\text{ and all }c_{j}\in N(c_{i}),$
$P_{\boldsymbol{\gamma}}({\mathbf{B}}_{c_{\boldsymbol{\gamma}}})=\boldsymbol{\mu}_{\boldsymbol{\gamma}},\quad\text{for }c_{\boldsymbol{\gamma}}\in{\mathcal{Q}}_{\Gamma}{\subseteq{\mathcal{Q}}_{S}},\qquad{\langle}{\mathbf{B}}_{c},\mathbf{1}{\rangle}=1,\quad\text{for }c{\in{\mathcal{Q}}_{S}\backslash{\mathcal{Q}}_{\Gamma}.}$
The separators in ${\mathcal{Q}}_{S}$ enforce that the cost cliques are consistent with each other, and thereby play a similar role as the separator sets in Definition 1. To develop an algorithm for problem 42 we add a local regularization term for each clique that can contain a cost, i.e., for each $c\in{\mathcal{Q}}_{C}$. Therefore, we generalize the expression for the reference joint distribution in 29. For $c\in{\mathcal{Q}}_{C}$, let $c_{j},c_{k}\in{\mathcal{Q}}_{S}$ be the two neighbours, and define ${\bf M}_{c}=\bm{}{c_{j}}\otimes\bm{}{c_{k}}$, where $\bm{}{c_{i}}=\boldsymbol{\mu}_{\boldsymbol{\gamma}}\otimes\mathbf{1}$ if ${\boldsymbol{\gamma}}\subset c_{i}$ for some ${\boldsymbol{\gamma}}\subset\Gamma$, and otherwise $\bm{}{c_{i}}=\mathbf{1}$. We use the notation $\otimes\mathbf{1}$ to broadcast a tensor to a bigger tensor of appropriate size. In the following, the broadcasting dimensions will always be obvious from the context. For example, if ${c_{i}}=\{x_{1},x_{2},x_{3}\},{\boldsymbol{\gamma}}=\{x_{2}\},$ then $\bm{}{c_{i}}(x_{1},x_{2},x_{3})=\boldsymbol{\mu}_{\boldsymbol{\gamma}}(x_{2})$.
Moreover, we add a mass constraint similar to 30. The regularized problem thus reads
$\min_{{\mathbf{B}}_{c},c\in{\mathcal{Q}}}$ $\sum_{c\in{\mathcal{Q}}_{C}}{\langle}{\bf C}_{c},{\mathbf{B}}_{c}{\rangle}+\epsilon{\mathcal{H}}({\mathbf{B}}_{c}|{\bf M}_{c})$ (43)
subject to $P_{{c_{i}}\cap{c_{j}}}({\mathbf{B}}_{c_{i}})=P_{{c_{i}}\cap{c_{j}}}({\mathbf{B}}_{c_{j}}),\quad\text{ for each }c_{i}\in{\mathcal{Q}}_{S},\text{ and all }c_{j}\in N(c_{i})$
$P_{\boldsymbol{\gamma}}({\mathbf{B}}_{c_{\boldsymbol{\gamma}}})=\boldsymbol{\mu}_{\boldsymbol{\gamma}},\quad\text{for }c_{\boldsymbol{\gamma}}\in{\mathcal{Q}}_{\Gamma}\subseteq{\mathcal{Q}}_{S},\qquad{\langle}{\mathbf{B}}_{c},\mathbf{1}{\rangle}=1,\quad\text{for }c\in{\mathcal{Q}}_{S}\backslash{\mathcal{Q}}_{\Gamma}.$
5.2 Algorithm
The problem in 43 is similar to 30, and we will extend the algorithm developed in Section 4 to handle general graphs. More precisely, following the derivations in Sec. 4, we can derive a method for approximating the solution of 43 similar to Algorithm 3. Note that the second set of constraints in 43 correspond to the constraints on $\Gamma$ in the tree-structured optimal transport problem 30. Special care needs to be applied to the first set of constraints. In particular, in order to derive an efficient algorithm we need the neighbour sets in Definition 4.4 to be nested as described in the following assumption.
Assumption 1.
For all $c\in{\mathcal{Q}}_{S}$ with at least two neighbors, there is a inclusion sequence $(c\cap c_{j_{1}})=(c\cap c_{j_{2}})\supseteq\dots\supseteq(c\cap c_{j_{\ell}})$ that includes all $\ell$ neighbours $c_{j_{1}},\dots,c_{j_{\ell}}$ of $c$. In case a constraint set ${\boldsymbol{\gamma}}\subset\Gamma$ is embedded in the clique $c$, then this can be added after the smallest set in the inclusion, i.e., $(c\cap c_{j_{\ell}})\supseteq{\boldsymbol{\gamma}}$.
Example 4.
We give an example of the nested structure in Assumption 1 in Fig. 9. Here, $c=\{x_{1},x_{3},x_{4}\},c_{1}=\{x_{1},x_{2},x_{3},x_{4}\},c_{2}=\{x_{1},x_{3},x_{4}\},c_{3}=\{x_{0},x_{1},x_{3}\}$, and finally ${\boldsymbol{\gamma}}=\{x_{3}\}.$
Figure 9: Example on nested structure.
1,2,3,412341,2,3,41 , 2 , 3 , 43333111144441,3,41341,3,41 , 3 , 40,1,30130,1,30 , 1 , 3
Proposition 2.
The solution to 43 is of the form ${\mathbf{B}}_{c}={\bf K}_{c}\odot{\bf U}_{c}\odot{\bf M}_{c},$ for each $c\in{\mathcal{Q}}_{C}$, where ${\bf K}_{c}=\exp(-{\bf C}_{c}/\epsilon)$ and ${\bf U}_{c}=\left({\mathbf{u}}^{c_{j}}_{i_{j}}\otimes\mathbf{1}\right)\odot\left({\mathbf{u}}^{c_{k}}_{i_{k}}\otimes\mathbf{1}\right)$ for two smaller tensors ${\mathbf{u}}^{c_{j}}_{i_{j}}$ and ${\mathbf{u}}^{c_{k}}_{i_{k}}$. These tensors are dual variables associated with the two neighbors $c_{j},c_{k}\in{\mathcal{Q}}_{S}$ of $c$, and $i_{j}$, $i_{k}$ denote the position in the corresponding inclusion sequence as defined in Assumption 1. More precisely, denoting $\ell_{c}$ the length of the inclusion sequence for each $c\in{\mathcal{Q}}_{S}$, the dual problem for (43) reads
$\max_{\begin{subarray}{c}{\mathbf{u}}^{c}_{i},i=1,\ldots,\ell_{c}\text{ for }c\in{\mathcal{Q}}_{S}\\ {\mathbf{u}}_{{\boldsymbol{\gamma}}}^{c}\text{ for }c_{\boldsymbol{\gamma}}\in{\mathcal{Q}}_{\Gamma}\\ \rho_{c}\text{ for }c\in{\mathcal{Q}}_{S}\backslash{\mathcal{Q}}_{\Gamma}\end{subarray}}$ $-\sum_{c\in{\mathcal{Q}}_{C}}\|{\bf K}_{c}\odot{\bf U}_{c}\odot{\bf M}_{c}\|_{1}+\sum_{c_{\boldsymbol{\gamma}}\in{\mathcal{Q}}_{\Gamma}}\langle\boldsymbol{\mu}_{\boldsymbol{\gamma}},\log({\mathbf{u}}^{c}_{{\boldsymbol{\gamma}}})\rangle+\sum_{c\in{\mathcal{Q}}_{S}\backslash{\mathcal{Q}}_{\Gamma}}\rho_{c}$ (44)
subject to $\sum_{i=1}^{\ell_{c}}\log({\mathbf{u}}^{c}_{i}\otimes\mathbf{1})=\log({\mathbf{u}}^{c}_{{\boldsymbol{\gamma}}}\otimes\mathbf{1}),\quad\text{for}c\in{\mathcal{Q}}_{\Gamma}$
$\sum_{i=1}^{\ell_{c}}\log({\mathbf{u}}^{c}_{i}\otimes\mathbf{1})=\rho_{c}\mathbf{1},\quad\text{for }c\in{\mathcal{Q}}_{S}\backslash{\mathcal{Q}}_{\Gamma}.$
The proof of proposition 2 follows the standard Lagrangue dual theory. Recall that any separator set $c\in{\mathcal{Q}}_{S}$ has a set of neighbours $c_{1},\dots,c_{\ell}\in{\mathcal{Q}}_{C}$ that define an inclusion sequence $(c\cap c_{1})=(c\cap c_{2})\supseteq\dots\supseteq(c\cap c_{\ell})$ as in Assumption 1. To find the optimal dual variables of the dual problem (44) we propose to iterate through all $c\in{\mathcal{Q}}_{S}$, and update the dual variables ${\mathbf{u}}^{c}_{1},\dots\mathbf{u}^{c}_{\ell_{c}}$.
Due to the fourth condition in Definition 4, the modified junction tree forms a bipartite graph, when we consider the separators in ${\mathcal{Q}}_{S}$ as nodes, and the cliques in ${\mathcal{Q}}_{C}$ as edges. For example, in Fig. 7, the square node $\{0,J+1\}$ alone is one set in a bipartite graph, and the bottom gray nodes belong to another set. The bipartite structure allows for a parallelized algorithm. In each update we can apply coordinate ascent to all the variables associated with separators in one set together. Moreover, because those variables do not depend on each other, we can consider a subproblem of the dual, as we will describe in the following. Note that the iterative Bregman projection (Benamou et al. 2015) applies a similar idea to the Wasserstein barycenter problem. To be precise, for the updates related to each $c\in{\mathcal{Q}}_{S}$ it suffices to consider a subproblem of the dual (44) , which for $c\in{\mathcal{Q}}_{\Gamma}$ can be formulated as
$\max_{\begin{subarray}{c}{\mathbf{u}}_{i},i=1,\ldots,\ell\\ {\mathbf{v}}_{\ell+1}\end{subarray}}\sum_{i=1}^{\ell}-{\langle}{\mathbf{k}}_{i},{\mathbf{u}}_{i}{\rangle}+\langle\boldsymbol{\mu}_{\boldsymbol{\gamma}},\log({\mathbf{v}}_{\ell+1})\rangle\qquad\mbox{subject to }\sum_{i=1}^{\ell}\log({\mathbf{u}}_{i}\otimes\mathbf{1})=\log({\mathbf{v}}_{\ell+1}\otimes\mathbf{1}).$ (45)
Here we have omitted the superscript of the dual variables. Moreover, we denote the dual variable ${\mathbf{u}}_{\gamma}={\mathbf{v}}_{\ell+1}$, the reason for which will become apparent in the algorithm. Also, we introduce the notation ${{\mathbf{k}}_{i}:=P_{c\cap c_{i}}({\bf K}_{c_{i}}\odot{\bf M}_{c_{i}}\odot(\bar{\mathbf{u}}_{i}\otimes\mathbf{1}))}$, where $\bar{\mathbf{u}}_{i}$ denotes the second dual variable defining the tensor ${\bf U}_{c_{i}}$ together with ${\mathbf{u}}_{i}$, as described in Proposition 2. Thus, we can write $\|{\bf K}_{c_{i}}\odot{\bf U}_{c_{i}}\odot{\bf M}_{c_{i}}\|_{1}={\langle}{\mathbf{k}}_{i},{\mathbf{u}}_{i}{\rangle}$. Note that the case $c\in{\mathcal{Q}}_{S}\backslash{\mathcal{Q}}_{\Gamma}$ is a special case of (45), where ${\mathbf{v}}_{\ell+1}={\exp(\rho_{c})}\in\mathbb{R}$ is a scalar and $\boldsymbol{\mu}_{\gamma}=1$.
The optimal dual variables of the dual problem (44) can now be found by iterating through all $c\in{\mathcal{Q}}_{S}$, and solving a problem of the form (45). This can be done by performing the steps described in the following.
Proposition 3.
The solution to (45), where $\ell>1$, can be found by performing the following steps.
  1. (1)
    Calculate geometric mean variables from the largest intersection cluster $c\cap c_{1}$ to the smallest $c\cap c_{\ell}$ as
    ${\mathbf{q}}_{2}=\left({\mathbf{k}}_{1}\odot{\mathbf{k}}_{2}\right)^{1/2};\qquad{\mathbf{q}}_{i}=\left(P_{c\cap c_{i}}({\mathbf{q}}_{i-1})^{i-1}\odot{\mathbf{k}}_{i}\right)^{1/i},\qquad i=3,\dots,\ell$ (46)
  2. (2)
    Calculate intermediate variables from ${\mathbf{a}}_{\ell},\ldots,{\mathbf{a}}_{1}$ and ${\mathbf{v}}_{\ell},\ldots,{\mathbf{v}}_{3}$ from the smallest intersection cluster $c\cap c_{\ell}$ to the largest cluster $c\cap c_{1}$ as
    ${\mathbf{a}}_{i}=({\mathbf{v}}_{i+1}{\otimes\mathbf{1}})^{1/i}\odot{\mathbf{q}}_{i},\qquad\text{~{}~{}for }i=2,\ldots,\ell$ (47)
    ${\mathbf{v}}_{\ell+1}=(\boldsymbol{\mu}_{\gamma}./P_{\gamma}({\mathbf{q}}_{\ell}))^{\ell},\qquad{\mathbf{v}}_{i}=({\mathbf{a}}_{i}./P_{c\cap c_{i}}({\mathbf{q}}_{i-1}))^{i-1},\quad\text{for }i=3,\ldots,\ell$
  3. (3)
    Update ${\mathbf{u}}_{1},\ldots,{\mathbf{u}}_{\ell}$ as
    ${\mathbf{u}}_{1}={\mathbf{a}}_{2}./{\mathbf{k}}_{1},\quad{\mathbf{u}}_{i}={\mathbf{a}}_{i}./{\mathbf{k}}_{i}~{}~{}\mbox{ for }i={2,\dots,\ell},$ (48)
Proposition 3 only discusses the case of $\ell>1$. In the case $\ell=1$, the constraint in (45) asserts that ${\mathbf{u}}_{1}={\mathbf{v}}_{2}$. Thus, the objective in (45) is optimized by ${\mathbf{u}}_{1}=\boldsymbol{\mu}_{\gamma}./{\mathbf{k}}_{1}$.
Remark 3.
The method can be generalized to the case where the cost cliques have a larger number of neighbors, i.e., point 3. in Definition 4 is generalized to: For all $c\in C_{C}$ it holds $|N(c)|\leq k$ for some $k\geq 2$. A parallelized algorithm for optimal transport problems on this type of junction tree can be set up by considering a $k$-partite partition, instead of the bipartite partition in the algorithm above.
Figure 10: Number of iterations for converging to the accuracy $\delta=0.2$. The left column shows the number of iterations as a function of $d$ when $|E|$ is fixed ($|E|=3$ in (a) and $|E|=12$ in (b)), and the right column vice versa with $d=10$.
(a) Wasserstein Barycenter
(b) Wasserstein least square
6 Numerical examples
We present numerical results for two types of MOT problems. We examine the Wasserstein barycenter problem as discussed in Fan et al. (2022a, Example 1) and structured according to the graph in Fan et al. (2022a, Figure 1), and the Wasserstein Least Squares (WLS) example detailed in Sec. 2.2.2. The original graph and the modified junction tree for the WLS example are depicted in Fig. 1 and Fig. 7, respectively. For the sake of completeness, we also show modified junction tree for the Wasserstein spline example in Fig. 8. Notably, the barycenter problem is tree-structured, while the WLS problem forms a graph with a tree-width of two.
For the barycenter problem, the cost matrices $C_{(j,k)}$ in 28 are set as the squared Euclidean distances. In the WLS problem, the cost matrices follow 18 with coefficient $\alpha=10$. The constrained marginal distributions $\{\mu_{k}\}_{k\in\Gamma}$, supported on a uniform grid of $d$ points between 0 and 1, are generated from the log-normal distribution and normalized to sum to one. We set the accuracy threshold $\delta=0.2$ for the unregularized MOT problem.
Our comparison involves Algorithm 2, which solves a standard MOT problem with global regularization, and Algorithm 3, which addresses coupled bi-marginal OT with local regularization. We employ a random update rule for the global regularization method, consistent with the algorithm complexity analysis in Fan et al. (2022a). Our focus is on the number of iterations, anticipating that the parallel mechanism introduced by the local regularization method could reduce them. Each experiment is repeated five times with different random seeds, and the average number of iterations is reported in Fig. 10. Theoretical complexity bounds for the barycenter problem are also illustrated as dashed lines, with bounds for global and local regularization derived from Fan et al. (2022a, Theorem 2) and Theorem 3, respectively.
From the observations in Fig. 10(a), it is evident that the experimental results are consistent with the theoretical bounds. Notably, local regularization demonstrates the potential to reduce the number of iterations by a factor of $|E|$, leading to a lower iteration count beyond a certain threshold $|E|$. Furthermore, both global and local regularization exhibit a logarithmic dependency on the dimension, with a $\log(d)$ relationship. In contrast, for the results presented in Fig. 10(b), we have not established theoretical bounds for local regularization within this paper; this constitutes an area for future investigation.
7 Conclusion
In this paper, we study the multi-marginal optimal transport (MOT) problem with graph-structured costs. We propose a novel approach to solve the MOT problem using local entropy regularization. Specifically, we reformulate the entire MOT problem as a sum of smaller MOT problems. When the underlying graph is a tree, these smaller problems involve dual variables that can be divided into two groups, forming a bipartite graph. We develop an efficient parallel algorithm that alternately updates these two groups of dual variables. A key advantage of this approach is its parallelizability, making it well-suited for GPU implementation. We also provide a rigorous complexity analysis of this approach and extend it to handle MOT problems associated with general graphs.
Appendix A Proofs of Theorems
Proof.
Proof of Theorem 1 We perform a block-coordinate descent in 32. First, minimizing 32 with respect to $u_{(j,k_{j})}$ for $j\in\Gamma$, while keeping the other variables fixed, yields the first set of updates. To minimize the objective with respect to $\rho_{j}$ and $u_{(j,k)}$ for $j\notin\Gamma$, relax the constraint with Lagrangian multiplier $\beta$, which gives the Lagrangian
$L(\{\rho_{j}\},\{u_{(j,k)}\},\{\beta_{j}\})$ $=\sum_{(j,k)\in E}\|B_{(j,k)}\|_{1}-\sum_{j\in\Gamma}\log(u_{(j,k_{j})})^{\top}\mu_{j}$ (49)
$-\sum_{j\in V\backslash\Gamma}\left(\rho_{j}+\beta_{j}^{\top}\left(\sum_{k\in N(j)}\log(u_{(j,k)})-\rho_{j}\mathbf{1}\right)\right)$ (50)
The Lagrangian is minimized with respect to $u_{(j,k)}$ for $j\notin\Gamma$ if
$\beta_{j}=u_{(j,k)}\odot(K_{(j,k)}(u_{(k,j)}\odot\gamma_{k})),\quad\text{ for }k\in N(j).$ (51)
Multiplying these relations for all $k\in N(j)$ yields
$\bigodot_{k\in N(j)}\beta_{j}=e^{\rho_{j}}\bigodot_{k\in N(j)}\left(K_{(j,k)}(u_{(k,j)}\odot\gamma_{k})\right).$
Denote $v_{j}=\left(\bigodot_{k\in N(j)}(K_{(j,k)}(u_{(k,j)}\odot\gamma_{k}))\right)^{1/|N(j)|},$ it follows that $\beta_{j}=e^{\rho_{j}/|N(j)|}v_{j},$ which together with 51 leads to the updates for $u_{(j,k)}$, $j\notin\Gamma$. Note also that the Lagrangian 49 is only bounded in $\rho_{j}$ if it holds $\beta_{j}^{\top}\mathbf{1}=1$. Thus, $e^{\rho_{j}/|N(j)|}v_{j}^{\top}{\bf 1}=1$ and the update for $\rho_{j}$ follows. ∎
Proof.
Proof of Theorem 2 Define $\Delta_{t}:=f(u^{t},\rho^{t})-f(u^{*},\rho^{*}).$ Then according to the Lemma 3, Lemma 4, there is $\Delta_{t}-\Delta_{t+1}\geq\frac{1}{22|E|}\left(\max\left\{\frac{\epsilon\Delta_{t}}{2C_{\infty}},\delta^{\prime}\right\}\right)^{2}.$ On the one hand, we have $\frac{\epsilon^{2}\Delta_{t+1}}{88|E|C_{\infty}^{2}}\leq\frac{\epsilon^{2}\Delta_{t}}{88|E|C_{\infty}^{2}}-\left(\frac{\epsilon^{2}\Delta_{t}}{88|E|C_{\infty}^{2}}\right)^{2}.$ Note that $0\leq a_{t+1}\leq a_{t}-a_{t}^{2}$ implies $0\leq a_{t}\leq 1$. This implies whenever $\Delta_{t+1}\geq 0$, we have $0\leq\frac{\epsilon^{2}\Delta_{t}}{88|E|C_{\infty}^{2}}\leq 1$ for any $t>1$. Also note that $0\leq a_{t+1}\leq a_{t}-a_{t}^{2}$ implies ${1}/{a_{t+1}}-{1}/{a_{t}}\geq 1$ and further implies $1{1}/{a_{t}}-{1}/{a_{1}}\geq t-$ 1. Thus,
$t\leq 1+\frac{88|E|C_{\infty}^{2}}{\epsilon^{2}\Delta_{t}}-\frac{88|E|C_{\infty}^{2}}{\epsilon^{2}\Delta_{1}}.$ (52)
On the other hand,
$\Delta_{t+\tau}\leq\Delta_{t}-\frac{\delta^{\prime 2}\tau}{22|E|},\quad t,\tau\geq 0$ (53)
Fixing an number $\phi>0$, and considering using 52 for $\Delta_{t}\geq\phi$ and using 53 for $\Delta_{t}<\phi$. We minimize the sum of two estimates over the parameter $\phi\in(0,\Delta_{1}]$. Combining the above two inequalities, we get the total number of iterations satisfies
$t\leq\min_{0\leq\phi\leq\Delta_{1}}\left(1+\frac{88|E|C_{\infty}^{2}}{\epsilon^{2}\phi}-\frac{88|E|C_{\infty}^{2}}{\epsilon^{2}\Delta_{1}}+1+\frac{22|E|\phi}{\delta^{\prime 2}}\right)\leq 2+\frac{88|E|C_{\infty}}{\epsilon\delta^{\prime}}.$ (54)
Proof.
Proof of Theorem 3 With the specific choices $\epsilon=\frac{\delta}{4|E|\log(d)}$ and $\delta^{\prime}=\frac{\delta}{8C_{\infty}}$, Lemma 6 gives
$\sum_{(j,k)\in E}{\langle}C_{(j,k)},\widehat{B}_{(j,k)}{\rangle}-\sum_{(j,k)\in E}{\langle}C_{(j,k)},B^{*}_{(j,k)}{\rangle}\leq\delta.$ (55)
By Theorem 2, the stopping time $t$ satisfies $t\leq 2+\frac{88|E|C_{\infty}}{\delta^{\prime}\epsilon}=\mathcal{O}\left(\frac{|E|^{2}C_{\infty}^{2}\log(d)}{\delta^{2}}\right).$ Since each iteration of Algorithm 3 requires $\mathcal{O}(|E|d^{2})$ operations, a solution of Algorithm 3 is achieved in $\mathcal{O}\left(\frac{|E|^{3}C_{\infty}^{2}d^{2}\log(d)}{\delta^{2}}\right).$ On the other hand, Algorithm 4 takes $\mathcal{O}(|E|d^{2})$ (see Lemma 7 in Altschuler et al. (2017)). Therefore, the bound on total complexity follows. ∎
Appendix B Proofs of Propositions
Proof.
Proof of Proposition 1 Firstly, when $j\in\Gamma$, we have $\gamma_{j}=\mu_{j}$, and
[…]
When $j\notin\Gamma$, we have $\gamma_{j}={\bf 1}$. It follows that
[…]
$=\left(\bigodot_{k\in N(j)}q^{t}_{(j,k)})\right)^{{\frac{1}{|N(j)|}}}/e^{\rho_{j}^{t}/|N(j)|}=q^{t+1}_{j}/e^{\rho_{j}^{t}/|N(j)|},$ (58)
where we use the constraint $\sum_{k\in N(j)}\log(u_{(j,k)})=\rho_{j}\mathbf{1}$ of the dual problem 32 in the third equality, and the definition of $q_{j}^{t+1}$ in the last equality. This implies that
$\rho^{t+1}_{j}=-|N(j)|\log\left(\mathbf{1}^{\top}v^{t+1}_{j}\right)=\rho^{t}_{j}-|N(j)|\log({\bf 1}^{\top}q^{t+1}_{j})=\rho^{t}_{j}-|N(j)|\log(\|q^{t+1}_{j}\|_{1}).$ (59)
Furthermore,
[…]
Proof.
Proof of Proposition 3 We introduce auxiliary variables ${{\mathbf{v}}_{i}}$, $i=1,\dots,\ell$, and rewrite the constraints 45, resulting in the following reformulation of the problem
$\max_{\begin{subarray}{c}{\mathbf{u}}_{i},i=1,\ldots,\ell\\ {\mathbf{v}}_{i},i=3,\ldots,\ell+1\end{subarray}}$ $-\sum_{i=1}^{\ell}{\langle}{\mathbf{k}}_{i},{\mathbf{u}}_{i}{\rangle}+\langle\boldsymbol{\mu}_{\boldsymbol{\gamma}},\log({\mathbf{v}}_{\ell+1})\rangle$ (62)
subject to $\log({\mathbf{u}}_{1})+\log({\mathbf{u}}_{2})=\log({\mathbf{v}}_{3}\otimes\mathbf{1}),$ (63)
$\log({\mathbf{v}}_{i})+\log({\mathbf{u}}_{i})=\log({\mathbf{v}}_{i+1}\otimes\mathbf{1}),\quad\mbox{ for }i=3,\ldots,\ell-1$ (64)
$\log({\mathbf{v}}_{\ell})+\log({\mathbf{u}}_{\ell})=\log({\mathbf{v}}_{\ell+1}\otimes\mathbf{1}).$ (65)
Note here that ${\mathbf{v}}_{i}$ and ${\mathbf{u}}_{i}$ have the same dimension. The Lagrange relaxed problem of 62 is
$\max_{\begin{subarray}{c}{\mathbf{u}}_{i},i=1,\ldots,\ell\\ {\mathbf{v}}_{i},i=3,\ldots,\ell+1\end{subarray}}-\sum_{i=1}^{\ell}{\langle}{\mathbf{k}}_{i},{\mathbf{u}}_{i}{\rangle}+\langle\boldsymbol{\mu}_{\boldsymbol{\gamma}},\log({\mathbf{u}}_{\ell+1})\rangle+\langle{\mathbf{a}}_{2},\log({\mathbf{u}}_{1})+\log({\mathbf{u}}_{2})-\log({\mathbf{v}}_{3}\otimes\mathbf{1})\rangle$ (66)
$\qquad\qquad+\sum_{i=3}^{\ell-1}\langle{\mathbf{a}}_{i},\log({\mathbf{v}}_{i})+\log({\mathbf{u}}_{i})-\log({\mathbf{v}}_{i+1}\otimes\mathbf{1})\rangle+\langle{\mathbf{a}}_{\ell},\log({\mathbf{v}}_{\ell})+\log({\mathbf{u}}_{\ell})-\log({\mathbf{v}}_{\ell+1}\otimes\mathbf{1})\rangle,$
where ${\mathbf{a}}_{i}$ are newly introduced Lagrange variables. Maximizing with respect to ${\mathbf{v}}_{i}$ we get ${\mathbf{a}}_{i}=P_{c\cap c_{i}}({\mathbf{a}}_{i-1})$ for $i=3,\ldots,\ell$. Maximizing with respect to ${\mathbf{v}}_{\ell+1}$, we get that $P_{\boldsymbol{\gamma}}({\mathbf{a}}_{\ell})=\boldsymbol{\mu}_{\boldsymbol{\gamma}}.$ Further, maximizing with respect to ${\mathbf{u}}_{i}$, for $i=1,\dots,\ell$, we get 48. To satisfy the constraint (63) we need ${\mathbf{u}}_{1}\odot{\mathbf{u}}_{2}={\mathbf{v}}_{3}\otimes\mathbf{1}$. Plugging in the expressions for ${\mathbf{u}}_{1}$ and ${\mathbf{u}}_{2}$ in (48), we get the expression for ${\mathbf{a}}_{2}$ in (47) together with (46). The update of ${\mathbf{v}}_{3}$ in (47) follows from ${\mathbf{a}}_{3}=P_{c\cap c_{i}}({\mathbf{a}}_{2})$. Similarly, the remaining updates for ${\mathbf{q}}_{i}$, ${\mathbf{a}}_{i}$, and ${\mathbf{v}}_{i}$, for $i=3,\dots,\ell$, follow from the constraints ${\mathbf{v}}_{i}\odot{\mathbf{u}}_{i}={\mathbf{u}}_{\ell+1}\otimes\mathbf{1}$ and ${\mathbf{a}}_{i}=P_{c\cap c_{i}}({\mathbf{a}}_{i-1})$. Finally, the expression for ${\mathbf{v}}_{\ell+1}$ is a consequence of $\boldsymbol{\mu}_{\gamma}=P_{\gamma}({\mathbf{a}}_{\ell})$. ∎
Appendix C Proofs of Lemmas
Proof.
Proof of Lemma 1 When $j\in\Gamma\bigcap S$, the updates of $u_{(j,k)}$ would keep $\|B^{t+1}_{(j,k_{j})}\|_{1}=1$ since it matches the projection $B^{t+1}_{(j,k_{j})}\mathbf{1}$ and the true marginal $\mu_{j}$. When $j\in(V\backslash\Gamma)\bigcap S$, by proposition 1, the updates of $u_{(j,k)}$ are equivalent to 34 in Theorem 1, thus they satisfy 51, which lead to $\|B_{(j,k)}^{t+1}\|=\mathbf{1}^{\top}\beta^{t+1}_{j}.$ Since $\mathbf{1}^{\top}\beta^{t+1}_{j}=1$ holds, we get $\|B_{(j,k)}^{t+1}\|=1.$
Proof.
Proof of Lemma 2 By 33, for $j\in\Gamma$, following the logic in Dvurechensky et al. (2018, Lemma 1), there is
$\max_{i=1,\dots,d}(\lambda_{(j,k)}^{t})_{i}-\min_{i=1,\dots,d}(\lambda_{(j,k)}^{t})_{i}\leq\|C_{(j,k)}\|_{\infty}.$ (67)
By 34, for $j\in V\backslash\Gamma$, following Kroshnin et al. (2019, Lemma 4), there is
$\max_{i=1,\dots,d}(\lambda_{(j,k)}^{t})_{i}-\min_{i=1,\dots,d}(\lambda_{(j,k)}^{t})_{i}\leq\|C_{(j,k)}\|_{\infty}+\frac{1}{|N(j)|}\sum_{i\in N(j)}\|C_{(j,i)}\|_{\infty}.$ (68)
Combining these results gives the bound. ∎
Proof.
Proof of Lemma 3 By the gradient inequality for convex functions it holds that
$~{}~{}~{}~{}f(u^{t},\rho^{t})-f(u^{*},\rho^{*})=f(\lambda^{t},\rho^{t})-f(\lambda^{*},\rho^{*})$ (69)
$\leq\langle\nabla f(\lambda^{t},\rho^{t}),\left((\lambda^{t},\rho^{t})-(\lambda^{*},\rho^{*})\right)^{\top}\rangle$
$=\sum_{j\in\Gamma}\langle\lambda_{(j,k_{j})}^{t}-\lambda_{(j,k_{j})}^{*},\frac{1}{\epsilon}B_{(j,k_{j})}^{t}\mathbf{1}-\frac{1}{\epsilon}\mu_{j}\rangle$
$+\sum_{j\notin\Gamma}\sum_{k\in N(j)}\langle\lambda_{(j,k)}^{t}-\lambda_{(j,k)}^{*},\frac{1}{\epsilon}q_{(j,k)}^{t}\rangle-\sum_{j\notin\Gamma}(\rho_{j}^{t}-\rho_{j}^{*})$
$=\frac{1}{\epsilon}\sum_{j\in\Gamma}\langle\lambda_{(j,k_{j})}^{t}-\lambda_{(j,k_{j})}^{*},B_{(j,k_{j})}^{t}\mathbf{1}-\mu_{j}\rangle$
$+\sum_{j\notin\Gamma}\sum_{k\in N(j)}\left\langle\frac{1}{\epsilon}\lambda_{(j,k)}^{t}-\frac{1}{|N(j)|}\rho_{j}^{t}\mathbf{1}-\frac{1}{\epsilon}\lambda_{(j,k)}^{*}-\frac{1}{|N(j)|}\rho_{j}^{*}\mathbf{1}~{},~{}q_{(j,k)}^{t}\right\rangle.$
Here, we have used that $\langle\mathbf{1},q_{(j,k)}^{t}\rangle=1$, and thus $-\rho_{j}^{t}+\rho_{j}^{*}=\langle-\rho_{j}^{t}\mathbf{1}+\rho_{j}^{*}\mathbf{1},q_{(j,k)}^{t}\rangle$. We bound the first sum as in Fan et al. (2022a, proof of Lemma 2) (compare also Dvurechensky et al. (2018, proof of Lemma 2)), which leads to
$\sum_{j\in\Gamma}\langle\lambda_{(j,k_{j})}^{t}-\lambda_{(j,k_{j})}^{*},B_{(j,k_{j})}^{t}\mathbf{1}-\mu_{j}\rangle\leq 2C_{\infty}\sum_{j\in\Gamma}\|B_{(j,k_{j})}^{t}\mathbf{1}-\mu_{j}\|_{1}.$ (70)
To bound the second sum in the final expression of 69 note that for any $j\notin\Gamma,$
$\sum_{k\in N(j)}\left\langle\frac{1}{\epsilon}\lambda_{(j,k)}^{t}-\frac{1}{|N(j)|}\rho_{j}^{t}\mathbf{1}~{},~{}\bar{q}_{j}^{t}\right\rangle=\left\langle\sum_{k\in N(j)}\frac{1}{\epsilon}\lambda_{(j,k)}^{t}-\rho_{j}^{t}\mathbf{1}~{},~{}\bar{q}_{j}^{t}\right\rangle=0,$
since $\sum_{k\in N(j)}\lambda_{(j,k)}^{t}/\epsilon=\rho_{j}^{t}\mathbf{1}$. The same equality holds with $\lambda_{(j,k)}^{*}$ and $\rho_{j}^{*}$. Moreover, note that it holds
$\sum_{k\in N(j)}\langle\mathbf{1},q_{(j,k)}^{t}-\bar{q}_{j}^{t}\rangle=\left\langle\mathbf{1},\sum_{k\in N(j)}q_{(j,k)}^{t}-|N(j)|\bar{q}_{j}^{t}\right\rangle=0.$
Using these relations it follows
$~{}~{}~{}\sum_{k\in N(j)}\left\langle\frac{1}{\epsilon}\lambda_{(j,k)}^{t}-\frac{1}{|N(j)|}\rho_{j}^{t}\mathbf{1}-\frac{1}{\epsilon}\lambda_{(j,k)}^{*}-\frac{1}{|N(j)|}\rho_{j}^{*}\mathbf{1}~{},~{}q_{(j,k)}^{t}\right\rangle$ (71)
$=\frac{1}{\epsilon}\sum_{k\in N(j)}\left\langle\lambda_{(j,k)}^{t}-\lambda_{(j,k)}^{*}~{},~{}q_{(j,k)}^{t}-\bar{q}_{j}^{t}\right\rangle\leq\frac{2C_{\infty}}{\epsilon}\sum_{k\in N(j)}\|B_{(j,k)}^{t}\mathbf{1}-\bar{q}_{j}^{t}\|_{1}.$ (72)
Here, the last step follows with the same approach as for 70 (see also Kroshnin et al. (2019, proof of Lemma 5)). Finally, combining 69, 70 and 72 completes the proof. ∎
Proof.
Proof of Lemma 4 Using the relations in Algorithm 3 and Lemma 1, we get
$f(u^{t},\rho^{t})-f(u^{t+1},\rho^{t+1})=\sum_{j\in\Gamma}(\log u_{(j,k_{j})}^{t+1}-\log u_{(j,k_{j})}^{t})^{\top}\mu_{j}+\sum_{j\notin\Gamma}(\rho_{j}^{t+1}-\rho_{j}^{t})$ (73)
$=$ $\sum_{j\in\Gamma}\left({{\mathcal{H}}}\left(\mu_{j}|B_{(j,k_{j})}^{t}\mathbf{1}\right)+1\right)-\sum_{j\notin\Gamma}|N(j)|\log\left(\|q_{j}^{t+1}\|_{1}\right).$ (74)
By Pinsker’s inequality the first sum can be bounded as
$\sum_{j\in\Gamma}\left({{\mathcal{H}}}\left(\mu_{j}|B_{(j,k_{j})}^{t}\mathbf{1}\right)+1\right)\geq\frac{1}{2}\sum_{j\in\Gamma}\|B_{(j,k_{j})}^{t}\mathbf{1}-\mu_{j}\|_{1}^{2}.$
Moreover, since $\|q_{j}^{t+1}\|_{1}\leq 1$ and $\|q_{(j,k)}^{t}\|_{1}=1$, we can bound
$-\log\left(\|q_{j}^{t+1}\|_{1}\right)\geq 1-\|q_{j}^{t+1}\|_{1}=\frac{1}{|N(j)|}\sum_{k\in N(j)}\|q_{(j,k)}^{t}\|_{1}-\|q_{j}^{t+1}\|_{1}$ (75)
$=$ $\mathbf{1}^{\top}\left(\frac{1}{|N(j)|}\sum_{k\in N(j)}q_{(j,k)}^{t}-q_{j}^{t+1}\right)\geq\frac{1}{11|N(j)|}\sum_{k\in N(j)}\|q_{(j,k)}^{t}-\bar{q}_{j}^{t}\|_{1}^{2},$ (76)
where the last inequality is proved in (Kroshnin et al. 2019, proof of Lemma 6). Therefore,
$f(u^{t},\rho^{t})-f(u^{t+1},\rho^{t+1})$ $\geq\frac{1}{2}\sum_{j\in\Gamma}\|B_{(j,k_{j})}^{t}\mathbf{1}-\mu_{j}\|_{1}^{2}+\frac{1}{11}\sum_{j\notin\Gamma}\sum_{k\in N(j)}\|q_{(j,k)}^{t}-\bar{q}_{j}^{t}\|_{1}^{2}$ (77)
$\geq\frac{1}{11}\left(\sum_{j\in\Gamma}\|B_{(j,k_{j})}^{t}\mathbf{1}-\mu_{j}\|_{1}^{2}+\sum_{j\notin\Gamma}\sum_{k\in N(j)}\|q_{(j,k)}^{t}-\bar{q}_{j}^{t}\|_{1}^{2}\right)$ (78)
$\geq\frac{1}{22|E|}\left(\sum_{j\in\Gamma}\|B_{(j,k_{j})}^{t}\mathbf{1}-\mu_{j}\|_{1}+\sum_{j\notin\Gamma}\sum_{k\in N(j)}\|q_{(j,k)}^{t}-\bar{q}_{j}^{t}\|_{1}\right)^{2},$ (79)
where we use Cauchy-Schwarz in the final inequality. ∎
Proof.
Proof of Lemma 5 Denote the partition where the nodes have mismatch as $S$. Without loss of generality, we assume $j\in S$ and $k\in V\backslash S$.
$\sum_{(j,k)\in E}{\langle}C_{(j,k)},B_{(j,k)}{\rangle}-\sum_{(j,k)\in E}{\langle}C_{(j,k)},\widehat{B}_{(j,k)}{\rangle}=\sum_{(j,k)\in E}{\langle}C_{(j,k)},B_{(j,k)}-\widehat{B}_{(j,k)}{\rangle}$ (80)
$=\sum_{j\in\Gamma}{\langle}C_{(j,k_{j})},B_{(j,k_{j})}-\widehat{B}_{(j,k_{j})}{\rangle}+\sum_{j\notin\Gamma}\sum_{k\in N(j)}{\langle}C_{(j,k)},B_{(j,k)}-\widehat{B}_{(j,k)}{\rangle}.$ (81)
[…]
$\sum_{j\in\Gamma}{\langle}C_{(j,k_{j})},B_{(j,k_{j})}-\widehat{B}_{(j,k_{j})}{\rangle}\leq\sum_{j\in\Gamma}\|C_{(j,k_{j})}\|_{\infty}\|B_{(j,k_{j})}-\widehat{B}_{(j,k_{j})}\|_{1}\leq 2C_{\infty}\sum_{j\in\Gamma}\|\mu_{j}-B_{(j,k)}{\bf 1}\|_{1}.$ (82)
Similarly, for the second part,
$~{}~{}~{}~{}\sum_{j\notin\Gamma}\sum_{k\in N(j)}{\langle}C_{(j,k)},B_{(j,k)}-\widehat{B}_{(j,k)}{\rangle}\leq\sum_{j\notin\Gamma}\sum_{k\in N(j)}\|C_{(j,k)}\|_{\infty}\|B_{(j,k)}-\widehat{B}_{(j,k)}\|_{1}$ (83)
$\leq 2\sum_{j\notin\Gamma}\sum_{k\in N(j)}\|C_{(j,k)}\|_{\infty}\|\bar{q}_{j}-B_{(j,k)}{\bf 1}\|_{1}\leq 2C_{\infty}\sum_{j\notin\Gamma}\sum_{k\in N(j)}\|\bar{q}_{j}-B_{(j,k)}{\bf 1}\|_{1}.$ (84)
Summing them up gives
$\sum_{(j,k)\in E}{\langle}C_{(j,k)},B_{(j,k)}{\rangle}-\sum_{(j,k)\in E}{\langle}C_{(j,k)},\widehat{B}_{(j,k)}{\rangle}$ (85)
$\leq$ $2C_{\infty}\left(\sum_{j\in\Gamma}\|\mu_{j}-B_{(j,k_{j})}{\bf 1})\|_{1}+\sum_{j\notin\Gamma}\sum_{k\in N(j)}\|\bar{q}_{j}(B)-B_{(j,k)}{\bf 1}\|_{1}\right).$ (86)
It follows the exact same way to obtain the same bound for $\sum_{(j,k)\in E}{\langle}C_{(j,k)},\widehat{B}_{(j,k)}{\rangle}-\sum_{(j,k)\in E}{\langle}C_{(j,k)},B_{(j,k)}{\rangle}$. ∎
Proof.
Proof of Lemma 6 Let $\{\widehat{Y}_{(j,k)}\}$ denote the output from Algorithm 4 with inputs $\{B^{*}_{(j,k)}\}$ and $\{\widetilde{B}_{(j,k_{j})}{\bf 1}\}_{j\in\Gamma}$. Note that $\widetilde{B}_{(j,k)}$ is the optimal solution to
$\min_{B\in\Pi(\widetilde{B}_{(j,k)}{\bf 1},~{}\widetilde{B}^{\top}_{(j,k)}{\bf 1})}{\langle}C_{(j,k)},B{\rangle}+\epsilon{\mathcal{H}}(B|M_{(j,k)}),\quad\text{for }\forall(j,k)\in E$ (87)
which can be verified by checking KKT conditions. Thus it holds that
${\langle}C_{(j,k)},\widetilde{B}_{(j,k)}{\rangle}+\epsilon{\mathcal{H}}(\widetilde{B}_{(j,k)}|M_{(j,k)})\leq{\langle}C_{(j,k)},\widehat{Y}_{(j,k)}{\rangle}+\epsilon{\mathcal{H}}(\widehat{Y}_{(j,k)}|M_{(j,k)}).$ (88)
Since ${\langle}\widetilde{B}_{(j,k)},\log(\widetilde{B}_{(j,k)}){\rangle}\geq-2\log(d)$ and ${\langle}\widehat{Y}_{(j,k)},\log(\widehat{Y}_{(j,k)}){\rangle}\leq 0$ it follows that
$~{}~{}~{}~{}{\langle}C_{(j,k)},\widetilde{B}_{(j,k)}{\rangle}-{\langle}C_{(j,k)},\widehat{Y}_{(j,k)}{\rangle}\leq\epsilon{\mathcal{H}}(\widehat{Y}_{(j,k)}|M_{(j,k)})-\epsilon{\mathcal{H}}(\widetilde{B}_{(j,k)}|M_{(j,k)})$ (89)
$\leq-\epsilon{\langle}\widetilde{B}_{(j,k)},\log(\widetilde{B}_{(j,k)}){\rangle}+\epsilon{\langle}\widetilde{B}_{(j,k)}-\widehat{Y}_{(j,k)},\log M_{(j,k)}{\rangle}$ (90)
$\leq 2\epsilon\log(d)+\epsilon{\langle}\log\gamma_{j},\widetilde{B}_{(j,k)}{\bf 1}-\widehat{Y}_{(j,k)}{\bf 1}{\rangle}+\epsilon{\langle}\log\gamma_{k},\widetilde{B}_{(j,k)}^{\top}{\bf 1}-\widehat{Y}_{(j,k)}^{\top}{\bf 1}{\rangle}=2\epsilon\log(d),$ (91)
where the last equality holds because $\widetilde{B}_{(j,k)}{\bf 1}=\widehat{Y}_{(j,k)}{\bf 1}$ whenever $j\in\Gamma$, and $\gamma_{j}={\bf 1}$ whenever $j\notin\Gamma$.
The plans $\{B^{*}_{(j,k)}\}$ already satisfy the condition $\bar{q}_{j}(B^{*})=B^{*}_{(j,k)}{\bf 1}$ for $\forall j\in\Gamma,k\in N(j)$. Thus Lemma 5 gives
$\sum_{(j,k)\in E}{\langle}C_{(j,k)},\widehat{Y}_{(j,k)}{\rangle}-\sum_{(j,k)\in E}{\langle}C_{(j,k)},B^{*}_{(j,k)}{\rangle}$ (92)
$\leq 2C_{\infty}\left(\sum_{j\in\Gamma}\|\mu_{j}-\widetilde{B}_{(j,k_{j})}{\bf 1})\|_{1}\right),$ (93)
$\sum_{(j,k)\in E}{\langle}C_{(j,k)},\widehat{B}_{(j,k)}{\rangle}-\sum_{(j,k)\in E}{\langle}C_{(j,k)},\widetilde{B}_{(j,k)}{\rangle}$ (94)
$\leq 2C_{\infty}\left(\sum_{j\in\Gamma}\|\mu_{j}-\widetilde{B}_{(j,k_{j})}{\bf 1})\|_{1}+\sum_{j\notin\Gamma}\sum_{k\in N(j)}\|\bar{q}_{j}(\widetilde{B})-\widetilde{B}_{(j,k)}{\bf 1}\|_{1}\right).$ (95)
Summing up 91, 93 and 95 concludes the proof. ∎